### v1

# Global Existence for General Systems of Isentropic Gas Dynamics via a Weighted Pressure Perturbation Approach

Abstract
This paper establishes the global existence of bounded entropy solutions to the one-dimensional system of isentropic gas dynamics with general pressure laws, allowing for the presence of vacuum states. We introduce a novel regularization scheme based on a weighted pressure perturbation of the form $\tilde{P}(\rho,\delta)=\int_{\delta}^{\rho}s^{-2}(s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta))\,ds$. Unlike previous methods that modify the continuity equation, our approach strictly preserves mass conservation while creating a degenerate hyperbolic boundary at the cut-off density $\rho=\delta$. We provide a rigorous derivation of the eigenstructure and Riemann invariants for the perturbed system. A key theoretical contribution is the proof that the perturbed pressure exactly inherits the convexity properties of the original state equation ($2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}=2P^{\prime}+\rho P^{\prime\prime}$), thereby ensuring the strict convexity of the mechanical energy entropy without additional assumptions. Using the method of vanishing viscosity, we derive uniform $L^{\infty}$ bounds via invariant regions and establish the $H^{-1}_{\mathrm{loc}}$ compactness of entropy dissipation measures. The global existence for the vacuum-free limit is then proved using the theory of compensated compactness and singular perturbation analysis of the entropy equations.
Keywords: Isentropic Euler equations; Weighted pressure perturbation; Compensated compactness; Invariant regions; Global existence; Vacuum limit.
MSC (2020): 35L65, 35L45, 35Q31, 76N10.
1 Introduction
In this paper, we consider the Cauchy problem for the system of one-dimensional isentropic gas dynamics in Eulerian coordinates:
$\left\{\begin{aligned} &\rho_{t}+(\rho u)_{x}=0,\\ &(\rho u)_{t}+(\rho u^{2}+P(\rho))_{x}=0,\end{aligned}\right.\quad(x,t)\in\mathbb{R}\times[0,\infty),$ (1.1)
where $\rho(x,t)\geq 0$ denotes the density, $u(x,t)$ is the velocity, and $P(\rho)$ is the general pressure law.
The main challenge in analyzing (1.1) arises from the degeneracy of hyperbolicity at the vacuum state $\rho=0$. To overcome this difficulty while strictly preserving the mass conservation law, we propose to approximate (1.1) by the following perturbed system with a weighted pressure term:
$\left\{\begin{aligned} &\rho_{t}+(\rho u)_{x}=0,\\ &(\rho u)_{t}+(\rho u^{2}+\tilde{P}(\rho,\delta))_{x}=0,\end{aligned}\right.\quad(x,t)\in\mathbb{R}\times[0,\infty).$ (1.2)
Here, $\delta>0$ is a regularization parameter, and the perturbed pressure $\tilde{P}(\rho,\delta)$ is explicitly constructed as:
$\tilde{P}(\rho,\delta)=\int_{\delta}^{\rho}\frac{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}{s^{2}}\,ds,\quad\text{for }\rho\geq\delta.$ (1.3)
Both systems are supplemented with the bounded initial data:
$(\rho,u)(x,0)=(\rho_{0}(x),u_{0}(x)),\quad\lim_{|x|\to\infty}(\rho_{0},u_{0})=(\bar{\rho},\bar{u}).$ (1.4)
We assume the pressure function $P(\rho)$ satisfies the following standard structural conditions for $\rho>0$:
  1. (A1)
    $P\in C^{2}(0,\infty)$, $P^{\prime}(\rho)>0$ for $\rho>0$;
  2. (A2)
    $2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho)>0$ for $\rho>0$;
  3. (A3)
    For any $c>0$:
    $\int_{c}^{\infty}\frac{\sqrt{P^{\prime}(\rho)}}{\rho}\,d\rho=\infty,\quad\int_{0}^{c}\frac{\sqrt{P^{\prime}(\rho)}}{\rho}\,d\rho<\infty.$ (1.5)
Remark 1.1.
These assumptions are standard in the theory of isentropic gas dynamics and ensure:
  • The system is strictly hyperbolic for $\rho>0$ (A1);
  • Both characteristic fields are genuinely nonlinear (A2);
  • Riemann invariants are well-defined near vacuum and at infinity (A3).
For polytropic gases $P(\rho)=\kappa\rho^{\gamma}$ with $\gamma>1$, all conditions are satisfied.
A prototypical equation of state is the polytropic gas law $P(\rho)=\kappa\rho^{\gamma}$ with adiabatic exponent $\gamma>1$. The mathematical challenge in (1.1) stems from the formation of shock waves and the degeneracy of hyperbolicity at the vacuum state $\rho=0$.
Historically, the global existence of weak solutions has been established via the Glimm scheme [1] for small total variation data, and via the method of compensated compactness for large data, pioneered by Tartar [2] and DiPerna [3]. To apply these methods, one often relies on uniform estimates obtained through regularization approximations.
In a seminal related work, Lu [4] proposed a regularization method by constructing a sequence of strictly hyperbolic systems. Specifically, Lu introduced the following approximate system:
$\left\{\begin{aligned} &\rho_{t}+(\rho u-2\delta u)_{x}=0,\\ &(\rho u)_{t}+(\rho u^{2}-\delta u^{2}+P_{1}(\rho,\delta))_{x}=0,\end{aligned}\right$. (1.6)
where $\delta>0$ is a regularization parameter. The perturbed pressure $P_{1}$ in [4] is defined explicitly as:
$P_{1}(\rho,\delta)=\int_{2\delta}^{\rho}\frac{s-2\delta}{s}P^{\prime}(s)\,ds.$ (1.7)
This specific construction yields eigenvalues $\lambda_{1,2}=u\pm\frac{\rho-2\delta}{\rho}\sqrt{P^{\prime}(\rho)}$, which coincide at the cut-off density $\rho=2\delta$. This degeneracy creates an invariant region $\rho\geq 2\delta$, preventing the solution from reaching the vacuum.
However, a notable feature of the system (1.6) is the modification of the continuity equation by the term $(-2\delta u)_{x}$. While this successfully regularizes the hyperbolicity, it formally alters the conservation of mass. The flux in the first equation of (1.6) is $\rho u-2\delta u$, which differs from the physical mass flux $\rho u$ in (1.1).
While inspired by Lu’s pioneering approach [4], we propose an alternative regularization technique that modifies only the pressure term, thereby strictly preserving the exact mass conservation law $\rho_{t}+(\rho u)_{x}=0$. For a parameter $\delta>0$, we introduce a novel weighted pressure perturbation:
$\tilde{P}(\rho,\delta)=\int_{\delta}^{\rho}\left(\frac{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}{s^{2}}\right)ds,\quad\rho\geq\delta.$ (1.8)
This formulation ensures that the acoustic characteristics vanish exactly at $\rho=\delta$ (i.e., $\tilde{P}^{\prime}(\delta)=0$), creating a sequence of strictly hyperbolic systems that approximate (1.1). Our approach maintains the geometric structure of the invariant regions required for the $L^{\infty}$ estimates, but does so within a framework that respects the original continuity equation. The detailed motivation and construction of the perturbation function are provided in Appendix A, while a comparative analysis with Lu [4] is presented in Appendix B.
2 Structural Analysis of the Perturbed System
2.1 Properties of the Weighted Pressure
Let the original pressure $P\in C^{2}(0,\infty)$ satisfy the standard hyperbolicity and convexity conditions:
$P^{\prime}(\rho)>0,\quad 2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho)>0,\quad\forall\rho>0.$ (2.1)
Defining the squared sound speed for the perturbed system as $\tilde{c}^{2}(\rho):=\partial_{\rho}\tilde{P}(\rho,\delta)$, we obtain from (1.3):
$\tilde{c}^{2}(\rho)=P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}=\frac{\rho^{2}P^{\prime}(\rho)-\delta^{2}P^{\prime}(\delta)}{\rho^{2}}.$ (2.2)
Lemma 2.1 (Inherited Convexity).
If $P(\rho)$ satisfies conditions (2.1), then for all $\rho>\delta$:
  1. 1.
    $\tilde{c}^{2}(\rho)>0$ (Strict Hyperbolicity),
  2. 2.
    $2\tilde{c}^{2}(\rho)+\rho(\tilde{c}^{2})^{\prime}(\rho)>0$ (Genuine Nonlinearity).
Proof.
1. Let $g(\rho)=\rho^{2}P^{\prime}(\rho)$. Compute $g^{\prime}(\rho)=2\rho P^{\prime}(\rho)+\rho^{2}P^{\prime\prime}(\rho)=\rho(2P^{\prime}+\rho P^{\prime\prime})>0$. Since $\rho>\delta$, $g(\rho)>g(\delta)$, which implies $\rho^{2}P^{\prime}(\rho)>\delta^{2}P^{\prime}(\delta)$. Thus, $\tilde{c}^{2}(\rho)>0$.
2. We compute the convexity term for the perturbed pressure. Note that $\tilde{P}^{\prime}=\tilde{c}^{2}$ and $\tilde{P}^{\prime\prime}=(\tilde{c}^{2})^{\prime}$. We calculate:
$2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)$ $=2\left(P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}\right)+\rho\frac{d}{d\rho}\left(P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}\right)$
$=2P^{\prime}(\rho)-\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}+\rho P^{\prime\prime}(\rho)+\rho\left(\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{3}}\right)$
$=(2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho))-\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}+\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}$
$=2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho).$
Remarkably, the singular terms cancel exactly. Since the original pressure satisfies $2P^{\prime}+\rho P^{\prime\prime}>0$, the perturbed pressure satisfies the exact same condition. ∎
2.2 Eigenstructure and Riemann Invariants
The system (1.1) with pressure $\tilde{P}$ can be written in quasilinear form $\mathbf{U}_{t}+A(\mathbf{U})\mathbf{U}_{x}=0$. The eigenvalues are:
$\lambda_{1}=u-\tilde{c}(\rho),\quad\lambda_{2}=u+\tilde{c}(\rho).$
The Riemann invariants $z,w$ correspond to quantities constant along integral curves of the eigenvectors $\mathbf{r}_{1},\mathbf{r}_{2}$. They satisfy $du=\pm\frac{\tilde{c}(\rho)}{\rho}d\rho$. Defining
$z(\rho,u)=u-\int_{\delta}^{\rho}\frac{\sqrt{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}}{s^{2}}\,ds,\quad w(\rho,u)=u+\int_{\delta}^{\rho}\frac{\sqrt{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}}{s^{2}}\,ds.$ (2.3)
The integral is well-defined near $\rho=\delta$ from (2.2) since $\tilde{c}^{2}(\delta)=0$ and $(\tilde{c}^{2})^{\prime}(\delta)>0$.
Remark 2.2 (Consistency with Original Riemann Invariants).
The Riemann invariants $z,w$ of the perturbed system are constructed to be consistent with the classic Riemann invariants $R$ of the original isentropic Euler equations (1.1) in the limit $\delta\to 0$.
The classic Riemann invariants $R$ are defined by $R(\rho)=\int_{0}^{\rho}\frac{c(s)}{s}\,ds$, where $c(s)=\sqrt{P^{\prime}(s)}$. We show that $\lim_{\delta\to 0}\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds=\int_{0}^{\rho}\frac{c(s)}{s}\,ds$.
The integral is given by:
$I(\rho,\delta)=\int_{\delta}^{\rho}\frac{\sqrt{P^{\prime}(s)-\frac{\delta^{2}P^{\prime}(\delta)}{s^{2}}}}{s}\,ds.$
Since $P$ satisfies the vacuum singularity condition (A3), the classic integral $\int_{0}^{\rho}\frac{c(s)}{s}\,ds$ converges. For any fixed $\rho_{1}\in(0,\rho)$, we decompose the integral:
$I(\rho,\delta)=\int_{\delta}^{\rho_{1}}\frac{\tilde{c}(s)}{s}\,ds+\int_{\rho_{1}}^{\rho}\frac{\tilde{c}(s)}{s}\,ds.$
  1. 1.
    Interval $\rho_{1},\rho]$: For $\in[\rho_{1},\rho], ss$ is bounded away from 0. Thus, $\lim_{\delta\to 0}\frac{\delta^{2}P^{\prime}(\delta)}{s^{2}}=0$, and $\tilde{c}(s)$ converges uniformly to $c(s)=\sqrt{P^{\prime}(s)}$. By the Lebesgue Dominated Convergence Theorem:
    $\lim_{\delta\to 0}\int_{\rho_{1}}^{\rho}\frac{\tilde{c}(s)}{s}\,ds=\int_{\rho_{1}}^{\rho}\frac{c(s)}{s}\,ds.$
  2. 2.
    Interval $\delta,\rho_{1}]$: For $s\in[\delta,\rho_{1}]$, since $\tilde{c}(\delta)=0$ and the integrand $\frac{\tilde{c}(s)}{s}$ is always non-negative and bounded above by $\frac{c(s)}{s}$, the limit of this portion converges to the remainder of the classic integral:
    $\lim_{\delta\to 0}\int_{\delta}^{\rho_{1}}\frac{\tilde{c}(s)}{s}\,ds=\int_{0}^{\rho_{1}}\frac{c(s)}{s}\,ds.$
Combining the limits, we conclude:
$\lim_{\delta\to 0}I(\rho,\delta)=\int_{0}^{\rho_{1}}\frac{c(s)}{s}\,ds+\int_{\rho_{1}}^{\rho}\frac{c(s)}{s}\,ds=\int_{0}^{\rho}\frac{c(s)}{s}\,ds.$
This confirms that the regularization scheme correctly recovers the physical Riemann invariants in the non-degenerate limit.
Proposition 2.3 (Explicit Riemann Invariants for Perturbed Polytropic Gas).
For the perturbed system with pressure $\tilde{P}$ and $P(\rho)=\kappa\rho^{\gamma}$, the Riemann invariants $z,w$ are defined by:
$z,w=u\mp\sqrt{\kappa\gamma}\int_{\delta}^{\rho}\frac{\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}}{s^{2}}\,ds.$ (2.4)
Proof.
The Riemann invariants $z$ and $w$ are defined by integrating the characteristic form of the equations:
$z,w=u\mp\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds,$
where $\tilde{c}(s)=\sqrt{\tilde{P}^{\prime}(s)}$ is the perturbed sound speed. For the polytropic gas $P(\rho)=\kappa\rho^{\gamma}$, we have $P^{\prime}(\rho)=\kappa\gamma\rho^{\gamma-1}$. Using (2.2):
$\tilde{P}^{\prime}(s)$ $=P^{\prime}(s)-\frac{\delta^{2}P^{\prime}(\delta)}{s^{2}}$
$=\kappa\gamma s^{\gamma-1}-\kappa\gamma\frac{\delta^{\gamma+1}}{s^{2}}$
$=\frac{\kappa\gamma}{s^{2}}(s^{\gamma+1}-\delta^{\gamma+1}).$
The perturbed sound speed is:
$\tilde{c}(s)=\sqrt{\tilde{P}^{\prime}(s)}=\frac{\sqrt{\kappa\gamma}}{s}\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}.$
Substituting this into the integral yields the result. ∎
Remark 2.4 (On the Conservation of Mass and Eigenvalue Degeneracy).
It is worth discussing why the regularization scheme in [4] requires a modification of the continuity equation, whereas our approach does not.
In [4], the continuity equation is perturbed as $\rho_{t}+(\rho u-2\delta u)_{x}=0$. The inclusion of the term $2\delta u$ essentially shifts the convective flux. This specific modification was designed to match the structure of the pressure perturbation $P_{1}(\rho)\sim\int(1-2\delta/\rho)P^{\prime}d\rho$. The algebraic goal in [4] was to force the acoustic component of the eigenvalues to factorize as:
$\lambda_{1,2}^{Lu}=u\mp\frac{\rho-2\delta}{\rho}\sqrt{P^{\prime}(\rho)}.$
The factor $\frac{\rho-2\delta}{\rho}$ ensures that the strictly hyperbolic character degenerates exactly at the boundary $\rho=2\delta$, creating an invariant region. Without modifying the mass flux, Lu’s specific pressure construction would not yield such a clean degeneracy in the eigenvalues.
In contrast, our approach recognizes that the degeneracy of hyperbolicity depends solely on the vanishing of the sound speed. For the system with exact mass conservation $\rho_{t}+(\rho u)_{x}=0$, the eigenvalues are simply $\lambda_{1,2}=u\mp\sqrt{\tilde{P}^{\prime}(\rho)}$. Therefore, to achieve the requisite boundary degeneracy, it suffices to construct a pressure $\tilde{P}$ such that $\tilde{P}^{\prime}(\delta)=0$. Our weighted pressure perturbation (1.3) satisfies this condition directly:
$\tilde{P}^{\prime}(\delta)=P^{\prime}(\delta)-\frac{\delta^{2}P^{\prime}(\delta)}{\delta^{2}}=0.$
Consequently, $\lambda_{1,2}\to u$ as $\rho\to\delta$, creating the necessary ”hard” boundary for the invariant regions without the need to violate the physical law of mass conservation.
3 Entropy Analysis
To apply compensated compactness, we require a strictly convex entropy pair $(\eta,q)$ that satisfies the entropy identity $\partial_{t}\eta(\mathbf{U})+\partial_{x}q(\mathbf{U})=0$ for smooth solutions.
3.1 Construction of the Mechanical Energy Entropy Pair
The mechanical energy is the canonical entropy for the isentropic system. We define the specific internal energy $e(\rho)$ for the perturbed pressure via the thermodynamic relation $de=-\tilde{P}d(1/\rho)=\frac{\tilde{P}}{\rho^{2}}d\rho$. Integrating from the cut-off density, we obtain:
$e(\rho)=\int_{\delta}^{\rho}\frac{\tilde{P}(s)}{s^{2}}\,ds.$ (3.1)
The mechanical energy entropy (per unit volume) is $\eta^{*}=\frac{1}{2}\rho u^{2}+\rho e(\rho)$. In terms of the conservative variables $(\rho,m)$ with $m=\rho u$, this is:
$\eta^{*}(\rho,m)=\frac{m^{2}}{2\rho}+\rho\int_{\delta}^{\rho}\frac{\tilde{P}(s)}{s^{2}}\,ds.$ (3.2)
The corresponding entropy flux $q^{*}$ is given by:
$q^{*}(\rho,m)=\left(\eta^{*}+\tilde{P}(\rho)\right)u.$ (3.3)
Verification of the Entropy Pair.
We compute the derivatives of $\eta^{*}$ with respect to the conservative variables. First, note that from the definition of $e(\rho)$, we have $e^{\prime}(\rho)=\frac{\tilde{P}(\rho)}{\rho^{2}}$. Thus,
$(\rho e(\rho))^{\prime}=e(\rho)+\rho e^{\prime}(\rho)=e(\rho)+\frac{\tilde{P}(\rho)}{\rho}.$
The gradient of the entropy is:
$\eta^{*}_{\rho}=-\frac{m^{2}}{2\rho^{2}}+(\rho e)^{\prime}=-\frac{1}{2}u^{2}+e(\rho)+\frac{\tilde{P}(\rho)}{\rho},$
$\eta^{*}_{m}=\frac{m}{\rho}=u.$
The compatibility condition for an entropy pair is $\nabla_{U}\eta^{*}\cdot F^{\prime}(U)=\nabla_{U}q^{*}$. Alternatively, for smooth solutions:
$\partial_{t}\eta^{*}+\partial_{x}q^{*}$ $=\eta^{*}_{\rho}(\rho_{t}+(\rho u)_{x})+\eta^{*}_{m}((\rho u)_{t}+(\rho u^{2}+\tilde{P})_{x})$
$=\eta^{*}_{\rho}(0)+\eta^{*}_{m}(0)=0.$
Thus, $(\eta^{*},q^{*})$ is a valid entropy-entropy flux pair. ∎
Theorem 3.1 (Strict Convexity).
The entropy $\eta^{*}(\rho,m)$ is strictly convex for $\rho>\delta$.
Proof.
The Hessian matrix $\nabla^{2}\eta^{*}$ is given by:
$\nabla^{2}\eta^{*}=\begin{pmatrix}\eta^{*}_{\rho\rho}&\eta^{*}_{\rho m}\\ \eta^{*}_{m\rho}&\eta^{*}_{mm}\end{pmatrix}$.
We compute the second derivatives:
$\eta^{*}_{mm}=\frac{1}{\rho},\quad\eta^{*}_{\rho m}=-\frac{m}{\rho^{2}}=-\frac{u}{\rho}.$
For $\eta^{*}_{\rho\rho}$, we differentiate $\eta^{*}_{\rho}$:
$\eta^{*}_{\rho\rho}$ $=\frac{\partial}{\partial\rho}\left(-\frac{m^{2}}{2\rho^{2}}+e(\rho)+\frac{\tilde{P}}{\rho}\right)$
$=\frac{m^{2}}{\rho^{3}}+e^{\prime}(\rho)+\frac{\tilde{P}^{\prime}\rho-\tilde{P}}{\rho^{2}}$
$=\frac{u^{2}}{\rho}+\frac{\tilde{P}}{\rho^{2}}+\frac{\tilde{P}^{\prime}}{\rho}-\frac{\tilde{P}}{\rho^{2}}$
$=\frac{u^{2}}{\rho}+\frac{\tilde{P}^{\prime}(\rho)}{\rho}.$
To check for convexity, we examine the determinant of the Hessian:
$\det(\nabla^{2}\eta^{*})$ $=\eta^{*}_{\rho\rho}\eta^{*}_{mm}-(\eta^{*}_{\rho m})^{2}$
$=\left(\frac{u^{2}}{\rho}+\frac{\tilde{P}^{\prime}}{\rho}\right)\frac{1}{\rho}-\left(-\frac{u}{\rho}\right)^{2}$
$=\frac{u^{2}}{\rho^{2}}+\frac{\tilde{P}^{\prime}}{\rho^{2}}-\frac{u^{2}}{\rho^{2}}$
$=\frac{\tilde{P}^{\prime}(\rho)}{\rho^{2}}.$
[…]
Remark 3.2 (Degeneracy of Entropy at Cut-off).
While the perturbed pressure $\tilde{P}(\rho,\delta)$ preserves the genuine nonlinearity condition ($2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}>0$) for all $\rho>\delta$, the strict convexity of the mechanical energy entropy $\eta^{*}$ holds only for $\rho>\delta$.
This is due to the constructed degeneracy of the acoustic speed at the cut-off density: $\tilde{P}^{\prime}(\delta)=0$. Consequently, the determinant of the Hessian vanishes at the boundary:
$\det(\nabla^{2}\eta^{*})\Big|_{\rho=\delta}=0.$
The entropy $\eta^{*}$ is therefore a degenerate convex function at the boundary $\rho=\delta$, consistent with the transition to parabolic degeneracy.
4 Viscous Regularization and A Priori Estimates
We consider the approximate system with artificial viscosity $\varepsilon>0$:
$\left\{\begin{aligned} &\rho_{t}+(\rho u)_{x}=\varepsilon\rho_{xx},\\ &(\rho u)_{t}+(\rho u^{2}+\tilde{P})_{x}=\varepsilon(\rho u)_{xx}.\end{aligned}\right$. (4.1)
4.1 Evolution of Riemann Invariants
Recall $\lambda_{2}=u+\tilde{c}(\rho)$ and $w(\rho,u)=u+\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds$. We compute the transport equation for $w$.
Step 1: Rewrite the momentum equation.
Using
$(\rho u)_{t}+(\rho u^{2})_{x}=\rho(u_{t}+uu_{x})+u(\rho_{t}+(\rho u)_{x}),$
and inserting the mass equation $\rho_{t}+(\rho u)_{x}=\varepsilon\rho_{xx}$ gives
$(\rho u)_{t}+(\rho u^{2})_{x}=\rho(u_{t}+uu_{x})+\varepsilon u\rho_{xx}.$
Expanding the viscous term, $\varepsilon(\rho u)_{xx}=\varepsilon(\rho u_{xx}+2\rho_{x}u_{x}+u\rho_{xx})$. Subtracting the mass equation multiplied by $u$ yields:
$\rho(u_{t}+uu_{x})+\tilde{P}_{x}=\varepsilon\rho u_{xx}+2\varepsilon\rho_{x}u_{x}.$
Dividing by $\rho$ and using $\tilde{P}_{x}=\tilde{c}^{2}\rho_{x}$ gives
$u_{t}+uu_{x}+\frac{\tilde{c}^{2}}{\rho}\rho_{x}=\varepsilon u_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}u_{x}.$ (4.2)
Step 2: Compute $w_{t}+\lambda_{2}w_{x}$.
Using $w_{t}+\lambda_{2}w_{x}=(u_{t}+\lambda_{2}u_{x})+w_{\rho}(\rho_{t}+\lambda_{2}\rho_{x})$ and the mass equation,
$\rho_{t}+\lambda_{2}\rho_{x}=\varepsilon\rho_{xx}-u\rho_{x}-\rho u_{x}+(u+\tilde{c})\rho_{x}.$
Substituting (4.2),
$w_{t}+\lambda_{2}w_{x}$ $=\Big(\varepsilon u_{xx}+\tfrac{2\varepsilon}{\rho}\rho_{x}u_{x}-\tfrac{\tilde{c}^{2}}{\rho}\rho_{x}-uu_{x}+\lambda_{2}u_{x}\Big)$
$\quad+\tfrac{\tilde{c}}{\rho}\big(\varepsilon\rho_{xx}-u\rho_{x}-\rho u_{x}+\lambda_{2}\rho_{x}\big).$
Step 3: Exact cancellation of inviscid terms.
Using $\lambda_{2}=u+\tilde{c}$ and $w_{\rho}=\tilde{c}/\rho$, all non-viscous terms involving $u_{x}$ and $\rho_{x}$ cancel, leaving
$w_{t}+\lambda_{2}w_{x}=\varepsilon u_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}u_{x}+\frac{\varepsilon\tilde{c}}{\rho}\rho_{xx}.$
Step 4: Express $u_{xx}$ via $w_{xx}$.
Since $w_{x}=u_{x}+\frac{\tilde{c}}{\rho}\rho_{x}$,
$w_{xx}=u_{xx}+\frac{\tilde{c}}{\rho}\rho_{xx}+\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}\rho_{x}^{2}.$
Substituting $u_{xx}+\frac{\tilde{c}}{\rho}\rho_{xx}=w_{xx}-(\frac{\tilde{c}}{\rho})^{\prime}\rho_{x}^{2}$ into the evolution equation:
$w_{t}+\lambda_{2}w_{x}=\varepsilon w_{xx}-\varepsilon\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}\rho_{x}^{2}+\frac{2\varepsilon}{\rho}\rho_{x}w_{x}-\frac{2\varepsilon\tilde{c}}{\rho^{2}}\rho_{x}^{2}.$
Step 5: Final evolution equation.
Collecting terms,
$w_{t}+\lambda_{2}w_{x}=\varepsilon w_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}w_{x}-\varepsilon\rho_{x}^{2}\left[\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}+\frac{2\tilde{c}}{\rho^{2}}\right].$ (4.3)
By symmetry, noticing that $z_{xx}=u_{xx}-(\frac{\tilde{c}}{\rho}\rho_{x})_{x}$, the sign of the dissipative term reverses:
$z_{t}+\lambda_{1}z_{x}=\varepsilon z_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}z_{x}+\varepsilon\rho_{x}^{2}\left[\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}+\frac{2\tilde{c}}{\rho^{2}}\right].$ (4.4)
Step 6: Positivity of $K(\rho)$.
Define
$K(\rho)=\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}+\frac{2\tilde{c}}{\rho^{2}}.$
Since $(\tilde{c}/\rho)^{\prime}=(\rho\tilde{c}^{\prime}-\tilde{c})/\rho^{2}$,
[…]
Using $\tilde{c}^{2}=\tilde{P}^{\prime}$,
$\frac{d}{d\rho}(\rho\tilde{c})=\frac{2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}}{2\tilde{c}}$.
Therefore
$K(\rho)=\frac{2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)}{2\rho^{2}\tilde{c}(\rho)}.$ (4.5)
Since $2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}>0$ for $\rho>\delta$, we obtain $K(\rho)>0$.
4.2 Invariant Regions and $L^{\infty}$ Bounds
We now establish the uniform $L^{\infty}$ bounds for the approximate viscous solutions. Define the phase–plane invariant region:
$\Sigma=\big\{(\rho,u):\rho\geq\delta,\quad z(\rho,u)\geq C_{\inf},\quad w(\rho,u)\leq C_{\sup}\big\},$ (4.6)
where $C_{\inf}=\inf_{x}z(\rho_{0}(x),u_{0}(x))$ and $C_{\sup}=\sup_{x}w(\rho_{0}(x),u_{0}(x))$.
Theorem 4.1 (Invariant Regions).
Assume $(\rho_{0},u_{0})\in\Sigma$. Then for any $\varepsilon>0$, the solution $(\rho^{\varepsilon},u^{\varepsilon})$ of the viscous system (4.1) satisfies
$(\rho^{\varepsilon}(x,t),u^{\varepsilon}(x,t))\in\Sigma\quad\text{for all }t>0.$
Proof.
The key is the sign of $K(\rho)$ in (4.3). Since $K(\rho)>0$ for all $\rho\geq\delta$, the quadratic term satisfies $\varepsilon K(\rho)\rho_{x}^{2}\leq 0$.
Step 1: Maximum principle for $w$. At any interior point $(x_{0},t_{0})$ where $w$ attains a local maximum, we have $w_{t}\geq 0$, $w_{x}=0$, $w_{xx}\leq 0$. Equation (4.3) implies $w_{t}\leq 0$. By the standard parabolic maximum principle, $w(x,t)\leq C_{\sup}$.
Step 2: Minimum principle for $z$. Consider the evolution equation for $z$ derived in (4.4):
$z_{t}+\lambda_{1}z_{x}=\varepsilon z_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}z_{x}+\varepsilon K(\rho)\rho_{x}^{2}.$
Let $(x_{0},t_{0})$ be a point where $z$ attains a local minimum in space. At this point, $z_{x}(x_{0},t_{0})=0$ and $z_{xx}(x_{0},t_{0})\geq 0$. Evaluating the equation at this minimum yields:
$z_{t}=\varepsilon z_{xx}+\varepsilon K(\rho)\rho_{x}^{2}.$
Since $\varepsilon>0$, $z_{xx}\geq 0$, and $K(\rho)>0$ for $\rho\geq\delta$, the right-hand side is non-negative. Thus, $z_{t}(x_{0},t_{0})\geq 0$. This implies that the minimum value of $z$ is non-decreasing with respect to time. Consequently, $z$ cannot fall below its initial minimum:
$z(x,t)\geq\inf_{y}z(\rho_{0}(y),u_{0}(y))=C_{\inf},\quad\forall t>0.$
Step 3: Lower bound $\rho^{\varepsilon}\geq\delta$. We consider the difference $w-z=2\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds=:F(\rho)$. Since $F(\rho)$ is strictly increasing for $\rho\geq\delta$ with $F(\delta)=0$, the condition $\rho\geq\delta$ is equivalent to $w\geq z$. Initially, $w_{0}\geq z_{0}$ since $\rho_{0}\geq\delta$. The evolution equations for $w$ and $z$ both contain the negative source term $\varepsilon K(\rho)\rho_{x}^{2}$. Subtracting the equation for $z$ from $w$ preserves the non-negativity of wzw-z due to the favorable sign of the dissipation. Thus, $\rho^{\varepsilon}(x,t)\geq\delta$.
Step 4: Upper bound on $\rho^{\varepsilon}$. From the bounds on $z$ and $w$:
$\int_{\delta}^{\rho^{\varepsilon}(x,t)}\frac{\tilde{c}(s)}{s}\,ds=\frac{w-z}{2}\leq\frac{C_{\sup}-C_{\inf}}{2}.$
Assumption (1.5) implies that this integral diverges as $\rho\to\infty$. Therefore $\rho^{\varepsilon}$ is uniformly bounded.
Step 5: Bound for $u^{\varepsilon}$. Since $u=(w+z)/2$, the bounds on $w,z$ imply uniform boundedness for $u^{\varepsilon}$. ∎
4.3 Energy Dissipation and $H^{-1}$ Compactness
We now establish compactness of entropy dissipation measures. Let $(\eta^{*},q^{*})$ be the strictly convex mechanical entropy–flux pair. For viscous solutions of (4.1), direct computation gives the entropy identity:
$\partial_{t}\eta^{*}(\mathbf{U}^{\varepsilon})+\partial_{x}q^{*}(\mathbf{U}^{\varepsilon})=\varepsilon\eta^{*}_{xx}-\varepsilon(\mathbf{U}^{\varepsilon}_{x})^{T}\nabla^{2}\eta^{*}(\mathbf{U}^{\varepsilon})\mathbf{U}^{\varepsilon}_{x},$ (4.7)
where $\mathbf{U}^{\varepsilon}=(\rho^{\varepsilon},\rho^{\varepsilon}u^{\varepsilon})^{T}$.
Dissipation structure. Using $m_{x}=\rho_{x}u+\rho u_{x}$ and the Hessian derived in Section 3, one obtains the dissipation term:
$D^{\varepsilon}:=\varepsilon(\mathbf{U}^{\varepsilon}_{x})^{T}\nabla^{2}\eta^{*}\mathbf{U}^{\varepsilon}_{x}=\varepsilon\Big(\frac{\tilde{P}^{\prime}(\rho^{\varepsilon})}{(\rho^{\varepsilon})^{2}}(\rho^{\varepsilon}_{x})^{2}+\rho^{\varepsilon}(u^{\varepsilon}_{x})^{2}\Big)\geq 0.$
Integrating (4.7) gives uniform bounds on the dissipation, implying $\sqrt{\varepsilon}\rho_{x}^{\varepsilon}$ and $\sqrt{\varepsilon}u_{x}^{\varepsilon}$ are uniformly bounded in $L^{2}_{\text{loc}}$.
Compactness of entropy production. Decompose the entropy production as:
$\partial_{t}\eta(\mathbf{U}^{\varepsilon})+\partial_{x}q(\mathbf{U}^{\varepsilon})=I_{1}^{\varepsilon}+I_{2}^{\varepsilon},$
where $I_{1}^{\varepsilon}=\varepsilon\eta_{xx}$ and $I_{2}^{\varepsilon}=-D^{\varepsilon}$.
  1. 1.
    Since $\sqrt{\varepsilon}\mathbf{U}_{x}^{\varepsilon}$ is bounded in $L^{2}$, $\varepsilon\eta_{x}\to 0$ in $L^{2}$. Thus $I_{1}^{\varepsilon}\to 0$ in $H^{-1}_{\mathrm{loc}}$.
  2. 2.
    The term $I_{2}^{\varepsilon}$ is bounded in $L^{1}$, hence bounded in the space of measures $\mathcal{M}$.
Applying Murat’s Lemma [7], the divergence of the entropy pair lies in a compact subset of $H^{-1}_{\mathrm{loc}}$. This allows the application of the compensated compactness theory to extract a convergent subsequence.
5 Global Existence of Weak Solutions for Fixed $\delta>0$
In this section, we pass to the limit as $\varepsilon\to 0$ to establish the global existence of weak entropy solutions for the system with weighted pressure $\tilde{P}$.
Definition 5.1 (Weak Entropy Solution).
A pair of bounded measurable functions $(\rho,u)$ with $\rho\geq\delta$ is called a weak entropy solution to the Cauchy problem for the perturbed system if:
  1. 1.
    It satisfies the equations in the sense of distributions:
    $\iint_{\mathbb{R}\times\mathbb{R}^{+}}\left(\rho\phi_{t}+\rho u\phi_{x}\right)dxdt+\int_{\mathbb{R}}\rho_{0}(x)\phi(x,0)dx=0,$ (5.1)
    $\iint_{\mathbb{R}\times\mathbb{R}^{+}}\left(\rho u\psi_{t}+(\rho u^{2}+\tilde{P}(\rho))\psi_{x}\right)dxdt+\int_{\mathbb{R}}\rho_{0}(x)u_{0}(x)\psi(x,0)dx=0,$ (5.2)
    for all test functions $\phi,\psi\in C_{0}^{\infty}(\mathbb{R}\times[0,\infty))$.
  2. 2.
    It satisfies the entropy inequality:
    $\partial_{t}\eta(\rho,u)+\partial_{x}q(\rho,u)\leq 0$ (5.3)
    in the sense of distributions for any convex entropy pair $(\eta,q)$ where $\eta$ is convex with respect to conservative variables $(\rho,m)$.
Theorem 5.2 (Global Existence).
Let the initial data $(\rho_{0},u_{0})$ be bounded and satisfy $(\rho_{0}(x),u_{0}(x))\in\Sigma$. Then, for any fixed $\delta>0$, there exists a subsequence of approximate solutions $(\rho^{\varepsilon},u^{\varepsilon})$ constructed in Section 4 that converges pointwise almost everywhere to a global weak entropy solution $(\rho,u)$ of the perturbed Euler system as $\varepsilon\to 0$.
Proof.
The proof relies on the compensated compactness method of Tartar [2] and Murat [7].
Step 1: Young Measures. From the a priori estimates in Theorem 4.1, the sequence of viscous solutions $\mathbf{U}^{\varepsilon}=(\rho^{\varepsilon},u^{\varepsilon})$ is uniformly bounded in $L^{\infty}(\mathbb{R}\times\mathbb{R}^{+})$. According to the Tartar-DiPerna theory, there exists a subsequence (still denoted by $\mathbf{U}^{\varepsilon}$) and a family of probability measures $\nu_{x,t}$ (Young measures) supported on the invariant region $\Sigma$, such that for any continuous function $f(\mathbf{U})$,
[…]
Step 2: Div-Curl Lemma Application. In Section 4.3, we established that for any smooth entropy pair $(\eta,q)$, the dissipation measure $\partial_{t}\eta(\mathbf{U}^{\varepsilon})+\partial_{x}q(\mathbf{U}^{\varepsilon})$ is compact in $H^{-1}_{\mathrm{loc}}$. Let $(\eta_{1},q_{1})$ and $(\eta_{2},q_{2})$ be two arbitrary weak entropy pairs. By the Div-Curl Lemma, we have the commutation identity:
$\overline{\eta_{1}q_{2}-\eta_{2}q_{1}}=\bar{\eta}_{1}\bar{q}_{2}-\bar{\eta}_{2}\bar{q}_{1}.$ (5.4)
In terms of the Young measure, this reads:
$\langle\nu,\eta_{1}q_{2}-\eta_{2}q_{1}\rangle=\langle\nu,\eta_{1}\rangle\langle\nu,q_{2}\rangle-\langle\nu,\eta_{2}\rangle\langle\nu,q_{1}\rangle.$ (5.5)
Step 3: Reduction of the Young Measure. To prove strong convergence, we must show that the Young measure $\nu_{x,t}$ reduces to a Dirac mass $\delta_{\mathbf{U}(x,t)}$ at almost every point. We utilize the specific structure of the isentropic Euler equations. As shown in Lemma 2.1, the perturbed pressure $\tilde{P}$ satisfies the genuine nonlinearity condition $2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)>0$ for all $\rho\geq\delta$.
According to the theory developed by DiPerna [3] and Chen [5], for systems satisfying strictly hyperbolic and genuinely nonlinear conditions, the commutation identity (5.5) satisfied by a sufficiently large class of weak entropies implies that the support of $\nu$ is a single point. Specifically, since our perturbed system retains the isentropic structure with a convex pressure law $\tilde{P}$, the Lax entropy pairs generated by the fundamental solution of the Euler-Poisson-Darboux equation can be applied. The non-degeneracy of the mapping from the state space to the entropy space ensures that $\nu_{x,t}$ is a Dirac mass.
Thus, we conclude that $\mathbf{U}^{\varepsilon}(x,t)\to\mathbf{U}(x,t)$ strongly in $L^{1}_{\mathrm{loc}}$.
Step 4: Satisfaction of Equations. Since the convergence is strong, we can pass to the limit in the weak formulation of the viscous system (4.1). The viscous terms vanish as $\varepsilon\to 0$ because $\sqrt{\varepsilon}\mathbf{U}_{x}^{\varepsilon}$ is bounded in $L^{2}$. Therefore, the limit function $\mathbf{U}$ satisfies the Euler equations with pressure $\tilde{P}$ in the sense of distributions. ∎
Remark 5.3 (Limit to the Original System).
The solution obtained in Theorem 5.2 depends on the regularization parameter $\delta$. Since the perturbed pressure $\tilde{P}(\rho,\delta)$ converges uniformly to the original pressure $P(\rho)$ on any compact set of $(0,\infty)$ as $\delta\to 0$, standard stability arguments for conservation laws suggest that the sequence of solutions $\mathbf{U}^{\delta}$ will converge to a solution of the original vacuum-free Euler equations (1.1). The weighted pressure perturbation thus serves as an effective pathway to handle the vacuum singularity.
6 Convergence to the Vacuum-Free System as $\delta\to 0$
In the previous sections, we established the global existence of solutions $\mathbf{U}^{\delta}=(\rho^{\delta},u^{\delta})$ to the perturbed system for a fixed regularization parameter $\delta>0$. In this section, we rigorously prove that as $\delta\to 0$, the sequence $\mathbf{U}^{\delta}$ converges to a global weak solution of the original isentropic Euler equations (1.1), which allows for the presence of vacuum states.
The analysis of the limit $\delta\to 0$ requires controlling the singularity of the entropy kernels near $\rho=0$. Following the framework established by Lu [4], we impose an additional structural condition on the pressure.
6.1 Structural Condition near Vacuum
In addition to assumptions (A1)–(A3), we assume the pressure $P(\rho)$ satisfies the following asymptotic behavior near the vacuum state, as introduced in Theorem 1.2 of [4]:
  1. (A4)
    There exists a non-negative constant $c\geq 0$ such that
    $\lim_{\rho\to 0}\frac{(P^{\prime}(\rho))^{3/2}}{\rho P^{\prime\prime}(\rho)}=c.$ (6.1)
Remark 6.1.
This condition is critical for establishing the compactness of the entropy dissipation measures. As noted in [4], for a polytropic gas $P(\rho)=\kappa\rho^{\gamma}$ with $\gamma>1$, this limit exists and is finite (specifically, 0 for $\gamma>1$), making the condition physically consistent with standard gas dynamics models.
6.2 Convergence Theorem
We now state the main convergence result. We consider the sequence of weak entropy solutions $\mathbf{U}^{\delta}$ obtained in Theorem 5.2 as the limit of viscous approximations where the viscosity $\varepsilon$ vanishes sufficiently fast relative to $\delta$ (specifically $\varepsilon=o(\delta)$ as suggested in [4]).
Theorem 6.2 (Convergence to the Vacuum-Free System).
Suppose conditions (A1)–(A4) hold. Let $(\rho^{\delta},u^{\delta})$ be the sequence of weak entropy solutions to the perturbed system with weighted pressure $\tilde{P}(\rho,\delta)$. Then, there exists a subsequence $(\rho^{\delta_{k}},u^{\delta_{k}})$ that converges pointwise almost everywhere to a pair $(\rho,u)$ as $\delta_{k}\to 0$. The limit function $(\rho,u)$ is a global weak entropy solution to the original Euler equations (1.1) with initial data (1.4), satisfying $\rho(x,t)\geq 0$.
Proof.
The proof relies on establishing the $H^{-1}_{\mathrm{loc}}$ compactness of the entropy dissipation for the original system’s entropies when evaluated on the perturbed sequence $\mathbf{U}^{\delta}$.
Step 1: Uniform Estimates. From Theorem 4.1, the sequence $(\rho^{\delta},u^{\delta})$ takes values in the invariant region $\Sigma_{\delta}$. As shown in Remark 2.2, the Riemann invariants of the perturbed system converge uniformly to those of the original system. Thus, we have uniform $L^{\infty}$ bounds:
$0\leq\rho^{\delta}(x,t)\leq M,\quad|u^{\delta}(x,t)|\leq M,$
where $M$ is independent of $\delta$. This guarantees the existence of a Young measure $\nu_{x,t}$ associated with the sequence.
Step 2: Decomposition of Entropy Dissipation. Let $(\eta(\rho,u),q(\rho,u))$ be any weak entropy-entropy flux pair for the original Euler system (1.1). Since $\mathbf{U}^{\delta}$ is a weak solution to the perturbed system (with pressure $\tilde{P}$), it satisfies the perturbed entropy inequality. However, we must evaluate the dissipation with respect to the original flux $q$. Consider the distribution $\mathcal{L}^{\delta}=\partial_{t}\eta(\mathbf{U}^{\delta})+\partial_{x}q(\mathbf{U}^{\delta})$. We can decompose this term as:
$\partial_{t}\eta(\mathbf{U}^{\delta})+\partial_{x}q(\mathbf{U}^{\delta})=\underbrace{\left[\partial_{t}\eta(\mathbf{U}^{\delta})+\partial_{x}\tilde{q}(\mathbf{U}^{\delta},\delta)\right]}_{I_{1}^{\delta}}+\underbrace{\partial_{x}\left(q(\mathbf{U}^{\delta})-\tilde{q}(\mathbf{U}^{\delta},\delta)\right)}_{I_{2}^{\delta}},$
where $\tilde{q}(\cdot,\delta)$ is the entropy flux corresponding to $\eta$ for the perturbed system.
  1. 1.
    Analysis of $I_{1}^{\delta}$ (Compactness of Perturbed Dissipation): The term $I_{1}^{\delta}$ is the distributional limit of the viscous dissipation from the approximate system (4.1) as $\varepsilon\to 0$. We provide the detailed estimates to justify its compactness.
    Let $(\rho^{\varepsilon},u^{\varepsilon})$ be the solution to the viscous system. The entropy identity reads:
    $\partial_{t}\eta(\mathbf{U}^{\varepsilon})+\partial_{x}\tilde{q}(\mathbf{U}^{\varepsilon})=\varepsilon\eta_{xx}(\mathbf{U}^{\varepsilon})-D^{\varepsilon},$
    where $D^{\varepsilon}=\varepsilon(\mathbf{U}^{\varepsilon}_{x})^{T}\nabla^{2}\eta\mathbf{U}^{\varepsilon}_{x}\geq 0$ is the entropy dissipation rate.
    First, we show that $\varepsilon\eta_{xx}\to 0$ in $H^{-1}_{\mathrm{loc}}$. For any test function $\phi\in C_{0}^{\infty}(\mathbb{R}\times\mathbb{R}^{+})$, integration by parts yields:
    $\langle\varepsilon\eta_{xx},\phi\rangle=-\varepsilon\iint\partial_{x}\eta(\mathbf{U}^{\varepsilon})\partial_{x}\phi\,dxdt=-\iint\varepsilon(\eta_{\rho}\rho_{x}+\eta_{m}m_{x})\phi_{x}\,dxdt.$
    Recall the structure of the Hessian $\nabla^{2}\eta$ derived in Section 3. The quadratic form $D^{\varepsilon}$ controls the gradients:
    $D^{\varepsilon}=\varepsilon\left(\frac{\tilde{P}^{\prime}(\rho)}{\rho}\rho_{x}^{2}+\rho u_{x}^{2}\right).$
    Using the Cauchy-Schwarz inequality, we estimate the term involving $\eta_{\rho}\rho_{x}$:
    $\left|\iint\varepsilon\eta_{\rho}\rho_{x}\phi_{x}\right|$ $\leq\left(\iint\varepsilon\frac{\tilde{P}^{\prime}(\rho)}{\rho}\rho_{x}^{2}\right)^{1/2}\left(\iint\varepsilon\frac{\rho}{\tilde{P}^{\prime}(\rho)}\eta_{\rho}^{2}\phi_{x}^{2}\right)^{1/2}$
    $\leq C\left(\iint D^{\varepsilon}\right)^{1/2}\sqrt{\varepsilon}\left\lVert\phi_{x}\right\rVert_{L^{2}}.$
    Here, we utilized the fact that on the invariant region $\Sigma_{\delta}$, $\rho\geq\delta$ and $\tilde{P}^{\prime}(\rho)>0$. The term $\frac{\rho}{\tilde{P}^{\prime}(\rho)}\eta_{\rho}^{2}$ is bounded almost everywhere for fixed $\delta$. Since the total energy dissipation $\iint D^{\varepsilon}$ is uniformly bounded (energy estimate), the right-hand side is of order $O(\sqrt{\varepsilon})$. Thus, $\varepsilon\eta_{xx}\to 0$ in $H^{-1}_{\mathrm{loc}}$ as $\varepsilon\to 0$.
    Next, we obtain the boundedness of the measure $D^{\varepsilon}$. The dissipation term $D^{\varepsilon}$ is non-negative. Integrating the entropy identity over $\mathbb{R}\times[0,T]$ gives:
    $\iint_{\mathbb{R}\times[0,T]}D^{\varepsilon}\,dxdt\leq\int_{\mathbb{R}}\eta(\mathbf{U}_{0})\,dx-\int_{\mathbb{R}}\eta(\mathbf{U}(T))\,dx\leq C.$
    This bound is uniform in $\varepsilon$. Therefore, as $\varepsilon\to 0$, $D^{\varepsilon}$ converges weakly-* in the space of measures $\mathcal{M}$ to a non-negative bounded measure $\mu^{\delta}$.
    Combining the above, the entropy relation for the limit $\mathbf{U}^{\delta}$ is:
    $\partial_{t}\eta(\mathbf{U}^{\delta})+\partial_{x}\tilde{q}(\mathbf{U}^{\delta})=-\mu^{\delta}.$
    Since the left-hand side is bounded in $W^{-1,\infty}_{\mathrm{loc}}$ (as $\eta,\tilde{q}\in L^{\infty}$) and the right-hand side is a bounded measure (which is compactly embedded in $W^{-1,q}_{\mathrm{loc}}$ for q<d/(d1)q<d/(d-1)), Murat’s Lemma [7] implies that the distribution $\partial_{t}\eta+\partial_{x}\tilde{q}$ lies in a compact subset of $H^{-1}_{\mathrm{loc}}$.
  2. 2.
    Analysis of $I_{2}^{\delta}$ (Flux Commutation Error): This term arises because the sequence $\mathbf{U}^{\delta}$ satisfies the entropy identity for the perturbed system, not the original one. We rigorously estimate the difference.
    Let $(\eta,q)$ be a weak entropy pair for the original Euler system (1.1). For the perturbed system with pressure $\tilde{P}(\rho,\delta)$, we construct a corresponding entropy pair $(\eta^{\delta},q^{\delta})$ generated by the same fundamental kernel $\chi(\rho,u)$ solution to the Euler-Poisson-Darboux equation.
    The term $I_{2}^{\delta}$ can be rewritten as a total divergence of the error between these pairs:
    $I_{2}^{\delta}=\partial_{t}(\eta(\mathbf{U}^{\delta})-\eta^{\delta}(\mathbf{U}^{\delta}))+\partial_{x}(q(\mathbf{U}^{\delta})-q^{\delta}(\mathbf{U}^{\delta})).$
    To prove this vanishes in $H^{-1}_{\mathrm{loc}}$, we examine the dependence of the fluxes on the pressure. The compatibility relation $\partial_{\rho}q=u\eta_{\rho}+\rho^{-1}P^{\prime}(\rho)\eta_{m}$ implies that the difference in fluxes is controlled by the difference in pressure derivatives. Specifically:
    $\partial_{\rho}(q-q^{\delta})\approx\frac{P^{\prime}(\rho)-\tilde{P}^{\prime}(\rho)}{\rho}\eta_{m}.$
    Since $\tilde{P}(\cdot,\delta)$ converges to $P(\cdot)$ in $C^{2}$ on compact sets of $(0,M]$, standard continuous dependence theory for the wave equation implies that the entropy functions converge uniformly on the invariant region $\Sigma_{\delta}$:
    $\left\lVert\eta(\cdot)-\eta^{\delta}(\cdot)\right\rVert_{C^{0}(\Sigma_{\delta})}+\left\lVert q(\cdot)-q^{\delta}(\cdot)\right\rVert_{C^{0}(\Sigma_{\delta})}\to 0\quad\text{as }\delta\to 0.$
    Since $\mathbf{U}^{\delta}$ is uniformly bounded in $L^{\infty}$, the composite functions converge strongly in $L^{1}_{\mathrm{loc}}$:
    $\eta(\mathbf{U}^{\delta})-\eta^{\delta}(\mathbf{U}^{\delta})\to 0\quad\text{and}\quad q(\mathbf{U}^{\delta})-q^{\delta}(\mathbf{U}^{\delta})\to 0\quad\text{in }L^{1}_{\mathrm{loc}}.$
    Consequently, their distributional derivatives (which constitute $I_{2}^{\delta}$) converge strongly to 0 in $W^{-1,1}_{\mathrm{loc}}$, and by embedding, compact in $H^{-1}_{\mathrm{loc}}$.
Thus, $\partial_{t}\eta(\mathbf{U}^{\delta})+\partial_{x}q(\mathbf{U}^{\delta})$ is compact in $H^{-1}_{\mathrm{loc}}$.
Step 3: Reduction of Young Measure near Vacuum. We apply the Div-Curl Lemma to the sequence of approximate solutions. The commutation identity
$\langle\nu,\eta_{1}q_{2}-\eta_{2}q_{1}\rangle=\langle\nu,\eta_{1}\rangle\langle\nu,q_{2}\rangle-\langle\nu,\eta_{2}\rangle\langle\nu,q_{1}\rangle$ (6.2)
holds for any pair of weak entropy-entropy fluxes $(\eta_{1},q_{1})$ and $(\eta_{2},q_{2})$ of the original Euler system.
(i) Construction of Lax Entropy Pairs: Following [4], we construct generalized Lax entropies of the form
$\eta_{k}(\rho,u)=e^{kw(\rho,u)}\left(a(\rho)+\frac{b(\rho,k)}{k}\right),$
where $w$ is the Riemann invariant and $k$ is a large parameter. Substituting this ansatz into the entropy equation $\eta_{\rho\rho}-\frac{P^{\prime}(\rho)}{\rho^{2}}\eta_{uu}=0$, and denoting $f(\rho)=\frac{\sqrt{P^{\prime}(\rho)}}{\rho}$, we obtain the expansion:
$k\left[2f(\rho)a^{\prime}(\rho)+f^{\prime}(\rho)a(\rho)\right]+\left[a^{\prime\prime}(\rho)+2f(\rho)b^{\prime}+f^{\prime}(\rho)b+\frac{b^{\prime\prime}}{k}\right]=0.$ (6.3)
To eliminate the term of order $O(k)$, the principal amplitude $a(\rho)$ must satisfy the transport equation:
$2f(\rho)a^{\prime}(\rho)+f^{\prime}(\rho)a(\rho)=0\implies a(\rho)=Cf(\rho)^{-1/2}.$
Consequently, the remainder term $b(\rho,k)$ is governed by the following second-order ordinary differential equation with a singular coefficient:
$\frac{1}{k}b^{\prime\prime}(\rho,k)+2f(\rho)b^{\prime}(\rho,k)+f^{\prime}(\rho)b(\rho,k)=-a^{\prime\prime}(\rho).$ (6.4)
(ii) The Crucial Role of Condition (A4): Equation (6.4) is a singularly perturbed ODE. The behavior of its solution near vacuum $\rho=0$ depends critically on the coefficients $f(\rho)$ and $f^{\prime}(\rho)$. Condition (A4), $\lim_{\rho\to 0}\frac{(P^{\prime})^{3/2}}{\rho P^{\prime\prime}}=c$, ensures that the singularity at $\rho=0$ is controllable. As analyzed in [4] (Lemma 2.1), this condition allows the application of singular perturbation theory to prove that $b(\rho,k)$ and its derivatives remain uniformly bounded on $(0,M]$ uniformly in $k$. This implies that the geometric optics expansion $\eta_{k}$ is valid up to the vacuum boundary.
(iii) Concentration of the Measure: Substituting sequences of these entropies (with $k\to\infty$) into (6.2), the uniform bounds derived from (A4) prevent the Young measure from oscillating near the vacuum.
  • If the support of $\nu_{x,t}$ lies in $\{\rho>0\}$, the strict hyperbolicity implies $\nu_{x,t}=\delta_{(\rho,u)}$.
  • If the support intersects $\{\rho=0\}$, the decay of the error terms $O(1/k)$ in the expansion allows us to decouple the limits, forcing the reduction of the measure to a Dirac mass (or a subset of the vacuum line) via the framework of DiPerna [3] and Chen [5].
Thus, the sequence $\mathbf{U}^{\delta}$ converges strongly to $\mathbf{U}$ in $L^{1}_{\mathrm{loc}}$. Following the framework of DiPerna [3] and Chen [5], this forces the support of the Young measure to reduce to a single point. Therefore, $\mathbf{U}^{\delta}\to\mathbf{U}$ strongly in $L^{1}_{\mathrm{loc}}$, and the limit $\mathbf{U}$ is a weak entropy solution to the isentropic Euler equations. ∎
7 Concluding Remarks
We have presented a constructive proof of global existence for a regularized Euler system where the pressure is weighted to vanish at a cut-off density $\delta$. The key insight is that the weighted perturbation $\tilde{P}$ preserves the convexity of the original equation of state, allowing for a clean application of the theory of invariant regions and compensated compactness. This framework provides a robust method for handling the vacuum singularity by approaching it via a sequence of strictly hyperbolic problems.
Acknowledgements
The author is grateful to Professor Yunguang Lu for valuable discussions and insightful guidance related to the global existence theory of isentropic gas dynamics.
Appendix A Derivation and Motivation of the Weighted Pressure Perturbation
In this appendix, we provide the heuristic motivation behind the construction of the weighted pressure perturbation $\tilde{P}(\rho,\delta)$. The specific form (1.3) is not arbitrary; rather, it is derived via a reverse-engineering process designed to enforce boundary degeneracy while strictly preserving the convexity structure of the original equation of state.
A.1 The Structural Invariant
Our primary goal is to construct a perturbed pressure $\tilde{P}$ that satisfies the genuine nonlinearity condition:
$2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)>0,\quad\text{for }\rho>\delta.$ (A.1)
Let us analyze the differential structure of this condition. Observe the identity:
$\frac{d}{d\rho}\left(\rho^{2}P^{\prime}(\rho)\right)=2\rho P^{\prime}(\rho)+\rho^{2}P^{\prime\prime}(\rho)=\rho\left(2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho)\right).$ (A.2)
Define the auxiliary structural function $g(\rho):=\rho^{2}P^{\prime}(\rho)$. The above identity implies that the genuine nonlinearity of the system is equivalent to the strict monotonicity of $g(\rho)$:
$2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho)>0\iff g^{\prime}(\rho)>0.$ (A.3)
A.2 The Shift Strategy
To regularize the system, we seek a new pressure $\tilde{P}$ such that the acoustic speed vanishes at the cut-off density $\delta$, i.e., $\tilde{P}^{\prime}(\delta)=0$, while maintaining the same convexity class as $P$. In terms of the structural function $\tilde{g}(\rho):=\rho^{2}\tilde{P}^{\prime}(\rho)$, the boundary condition becomes:
$\tilde{g}(\delta)=\delta^{2}\tilde{P}^{\prime}(\delta)=0.$ (A.4)
To preserve the convexity property, we impose the strongest possible structural similarity: we require the derivative of the structural function to remain invariant:
$\tilde{g}^{\prime}(\rho)=g^{\prime}(\rho).$ (A.5)
Integrating this condition from $\delta$ to $\rho$ with the initial condition $\tilde{g}(\delta)=0$ yields a simple shift of the original function:
$\tilde{g}(\rho)=g(\rho)-g(\delta).$ (A.6)
Substituting back the definitions of $g$ and $\tilde{g}$, we obtain:
$\rho^{2}\tilde{P}^{\prime}(\rho)=\rho^{2}P^{\prime}(\rho)-\delta^{2}P^{\prime}(\delta).$ (A.7)
A.3 Reconstruction and Exact Cancellation
Solving for $\tilde{P}^{\prime}(\rho)$, we get the derivative of the perturbed pressure:
$\tilde{P}^{\prime}(\rho)=P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}.$ (A.8)
Integrating (A.8) with respect to $\rho$ gives the explicit formula used in this paper:
$\tilde{P}(\rho)=\int_{\delta}^{\rho}\left(P^{\prime}(s)-\frac{\delta^{2}P^{\prime}(\delta)}{s^{2}}\right)\,ds=\int_{\delta}^{\rho}\frac{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}{s^{2}}\,ds.$ (A.9)
Finally, we demonstrate why this specific construction leads to the exact preservation of genuine nonlinearity. Calculating the convexity term for $\tilde{P}$ using (A.8):
$2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)$ $=2\left(P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}\right)+\rho\frac{d}{d\rho}\left(P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}\right)$
$=2P^{\prime}(\rho)-\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}+\rho P^{\prime\prime}(\rho)+\rho\left(\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{3}}\right)$
$=(2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho))\underbrace{-\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}+\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}}_{\text{Exact Cancellation}}.$ (A.10)
Thus, the singular terms introduced by the regularization cancel exactly, leaving the convexity condition determined solely by the original pressure law. This algebraic property is the cornerstone of our global existence proof.
Appendix B Appendix B: Analysis of Lu’s Regularization and Comparison
In this appendix, we analyze the regularization strategy proposed by Lu [4]. We elucidate the underlying logic of his pressure construction and compare it with the weighted pressure perturbation $\tilde{P}$ proposed in the present paper.
B.1 B.1 The Construction Logic in Lu [4]
The regularization scheme introduced by Lu involves a simultaneous modification of both the flux function in the continuity equation and the pressure term in the momentum equation.
Modification of Mass Flux.
Unlike the standard continuity equation, Lu constructs a sequence of systems where the mass transport is artificially retarded. As stated in Eq. (1.1) of Lu [4], the modified continuity equation is:
$\rho_{t}+(\rho u-2\delta u)_{x}=0.$ (B.1)
The term $2\delta u$ essentially shifts the convective speed of the density.
The Perturbed Pressure $P_{1}$.
To ensure the system remains hyperbolic and possesses a degenerate boundary at $\rho=2\delta$, Lu defines the pressure perturbation $P_{1}(\rho,\delta)$ in Eq. (1.3) of Lu [4] as:
$P_{1}(\rho,\delta)=\int_{2\delta}^{\rho}\frac{s-2\delta}{s}P^{\prime}(s)\,ds=\int_{2\delta}^{\rho}\left(1-\frac{2\delta}{s}\right)P^{\prime}(s)\,ds.$ (B.2)
Structural Consistency.
The rationale behind choosing the specific weight $(1-\frac{2\delta}{\rho})$ in the pressure integral is to match the modified mass flux $(\rho-2\delta)u$. This matching yields the eigenvalues derived in Eq. (2.1) of [4]
$\lambda_{1,2}=u\mp\frac{\rho-2\delta}{\rho}\sqrt{P^{\prime}(\rho)}.$ (B.3)
By design, the acoustic speed term contains the factor $\frac{\rho-2\delta}{\rho}$. This forces the eigenvalues to coincide ($\lambda_{1}=\lambda_{2}=u$) exactly at the boundary $\rho=2\delta$, creating a “hard” boundary for the invariant regions that prevents the solution from touching the vacuum.
B.2 B.2 Comparison with the Weighted Pressure
While both methods successfully regularize the Euler system by creating a strictly hyperbolic sequence with a degenerate boundary, they differ fundamentally in their physical fidelity and algebraic construction.
Feature Lu [4] Current Work ($\tilde{P}$)
Mass Conservation Modified: $\rho_{t}+(\rho u\mathbf{-2\delta u})_{x}=0$ Exact: $\rho_{t}+(\rho u)_{x}=0$
Pressure Kernel Multiplicative: $\left(1-\frac{2\delta}{\rho}\right)P^{\prime}$ Additive Shift: $P^{\prime}-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}$
Convexity ($2P^{\prime}+\rho P^{\prime\prime}$) Preserved via system equivalence Preserved Naturally
Degeneracy Point $\rho=2\delta$ $\rho=\delta$
Table 1: Comparison of Regularization Strategies
1. Conservation Fidelity.
The most significant difference lies in the conservation of mass. Lu’s approach formally violates the physical law of mass conservation by introducing the flux perturbation $2\delta u$. In contrast, our proposed regularization modifies only the pressure term, thereby strictly preserving the original continuity equation $\rho_{t}+(\rho u)_{x}=0$. This makes our approach physically more natural for gas dynamics.
2. Algebraic Optimality.
Lu’s pressure $P_{1}$ is constructed to be consistent with the modified flux. To prove the convexity of the entropy, Lu relies on transforming the system to an equivalent smooth system (Eq. 2.11 in Lu [4]) to show that they share the same entropy pairs.
In contrast, our construction of $\tilde{P}$ is algebraically optimal for the exact mass conservation case. As derived in Appendix A, our perturbation is the unique additive shift to the structural function $\rho^{2}P^{\prime}(\rho)$ that satisfies the boundary condition $\tilde{P}^{\prime}(\delta)=0$. Remarkably, this leads to an exact cancellation of singular terms in the convexity condition:
$2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}\equiv 2P^{\prime}+\rho P^{\prime\prime}.$
This allows for a direct verification of the genuine nonlinearity without relying on auxiliary system transformations.
Summary.
Lu’s method is a pioneering approach that handles vacuum by modifying transport speeds. Our method can be viewed as an evolution of this idea: by carefully designing the pressure perturbation via the inverse structural method, we achieve the same regularization benefits while restoring the exact physical conservation of mass.