### v2

# Global Existence for General Systems of Isentropic Gas Dynamics via a Weighted Pressure Perturbation Approach

Abstract
This paper establishes the global existence of bounded entropy solutions to the one-dimensional system of isentropic gas dynamics with general pressure laws, allowing for the presence of vacuum states. We introduce a novel regularization scheme based on a weighted pressure perturbation designed to enforce acoustic degeneracy at a cut-off density $\rho=\delta$. A distinguishing feature of our approach is the strict preservation of the mass conservation law, in contrast to previous regularization methods that require modifying the continuity equation. We prove that the perturbed pressure $\tilde{P}$ exactly inherits the convexity structure of the original equation of state (satisfying $2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}=2P^{\prime}+\rho P^{\prime\prime}$), thereby ensuring the strict convexity of the mechanical entropy without auxiliary transformations. This structural preservation allows for the direct application of the Chueh-Conley-Smoller theory to establish uniform $L^{\infty}$ bounds via invariant regions. The convergence to a global weak solution in the vanishing regularization limit is finally proved using the method of compensated compactness and a singular analysis of the entropy equation near the vacuum.
Keywords: Isentropic Euler equations; Weighted pressure perturbation; Compensated compactness; Invariant regions; Global existence; Vacuum limit.
MSC (2020): 35L65, 35L45, 35Q31, 76N10.
1 Introduction
In this paper, we consider the Cauchy problem for the system of one-dimensional isentropic gas dynamics in Eulerian coordinates:
$\left\{\begin{aligned} &\rho_{t}+(\rho u)_{x}=0,\\ &(\rho u)_{t}+(\rho u^{2}+P(\rho))_{x}=0,\end{aligned}\right.\quad(x,t)\in\mathbb{R}\times[0,\infty),$ (1.1)
where $\rho(x,t)\geq 0$ denotes the density, $u(x,t)$ is the velocity, and $P(\rho)$ is the general pressure law.
A prototypical equation of state is the polytropic gas law $P(\rho)=\kappa\rho^{\gamma}$ with adiabatic exponent $\gamma>1$. The mathematical challenge in (1.1) stems from the formation of shock waves and the degeneracy of hyperbolicity at the vacuum state $\rho=0$.
Historically, the global existence of weak solutions has been established via the Glimm scheme [1] for small total variation data, and via the method of compensated compactness for large data, pioneered by Tartar [2] and DiPerna [3]. To apply these methods, one often relies on uniform estimates obtained through regularization approximations.
In a seminal related work, Lu [4] proposed a regularization method by constructing a sequence of strictly hyperbolic systems. Specifically, Lu introduced the following approximate system:
$\left\{\begin{aligned} &\rho_{t}+(\rho u-2\delta u)_{x}=0,\\ &(\rho u)_{t}+(\rho u^{2}-\delta u^{2}+P_{1}(\rho,\delta))_{x}=0,\end{aligned}\right$. (1.2)
where $\delta>0$ is a regularization parameter. The perturbed pressure $P_{1}$ in [4] is defined explicitly as:
$P_{1}(\rho,\delta)=\int_{2\delta}^{\rho}\frac{s-2\delta}{s}P^{\prime}(s)\,ds.$ (1.3)
This specific construction yields eigenvalues $\lambda_{1,2}=u\pm\frac{\rho-2\delta}{\rho}\sqrt{P^{\prime}(\rho)}$, which coincide at the cut-off density $\rho=2\delta$. This degeneracy creates an invariant region $\rho\geq 2\delta$, preventing the solution from reaching the vacuum.
However, a notable feature of the system (1.2) is the modification of the continuity equation by the term $(-2\delta u)_{x}$. While this successfully regularizes the hyperbolicity, it formally alters the conservation of mass. The flux in the first equation of (1.2) is $\rho u-2\delta u$, which differs from the physical mass flux $\rho u$ in (1.1).
While inspired by Lu’s pioneering approach [4], we propose an alternative regularization technique that modifies only the pressure term, thereby strictly preserving the exact mass conservation law $\rho_{t}+(\rho u)_{x}=0$. To overcome the vacuum difficulty, we approximate (1.1) by the following perturbed system with a novel weighted pressure perturbation:
$\left\{\begin{aligned} &\rho_{t}+(\rho u)_{x}=0,\\ &(\rho u)_{t}+(\rho u^{2}+\tilde{P}(\rho,\delta))_{x}=0,\end{aligned}\right.\quad(x,t)\in\mathbb{R}\times[0,\infty).$ (1.4)
Here, $\delta>0$ is a regularization parameter, and the perturbed pressure $\tilde{P}(\rho,\delta)$ is explicitly constructed as:
$\tilde{P}(\rho,\delta)=\int_{\delta}^{\rho}\frac{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}{s^{2}}\,ds,\quad\text{for }\rho\geq\delta.$ (1.5)
This formulation ensures that the acoustic characteristics vanish exactly at $\rho=\delta$ (i.e., $\tilde{P}^{\prime}(\delta)=0$), creating a sequence of strictly hyperbolic systems. Our approach maintains the geometric structure of the invariant regions required for the $L^{\infty}$ estimates, but does so within a framework that respects the original continuity equation. The detailed motivation and construction of the perturbation function are provided in Appendix A, while a comparative analysis with Lu [4] is presented in Appendix B.
Both systems are supplemented with the bounded initial data:
$(\rho,u)(x,0)=(\rho_{0}(x),u_{0}(x)),\quad\lim_{|x|\to\infty}(\rho_{0},u_{0})=(\bar{\rho},\bar{u}).$ (1.6)
We assume the pressure function $P(\rho)$ satisfies the following standard structural conditions for $\rho>0$:
  1. (A1)
    $P\in C^{2}(0,\infty)$, $P^{\prime}(\rho)>0$ for $\rho>0$;
  2. (A2)
    $2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho)>0$ for $\rho>0$;
  3. (A3)
    For any $c>0$:
    $\int_{c}^{\infty}\frac{\sqrt{P^{\prime}(\rho)}}{\rho}\,d\rho=\infty,\quad\int_{0}^{c}\frac{\sqrt{P^{\prime}(\rho)}}{\rho}\,d\rho<\infty.$ (1.7)
Remark 1.1.
These assumptions are standard in the theory of isentropic gas dynamics and ensure:
  • The system is strictly hyperbolic for $\rho>0$ (A1);
  • Both characteristic fields are genuinely nonlinear (A2);
  • Riemann invariants are well-defined near vacuum and at infinity (A3).
For polytropic gases $P(\rho)=\kappa\rho^{\gamma}$ with $\gamma>1$, all conditions are satisfied.
2 Structural Analysis of the Perturbed System
2.1 Properties of the Weighted Pressure
Let the original pressure $P\in C^{2}(0,\infty)$ satisfy the standard hyperbolicity and convexity conditions:
$P^{\prime}(\rho)>0,\quad 2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho)>0,\quad\forall\rho>0.$ (2.1)
Defining the squared sound speed for the perturbed system as $\tilde{c}^{2}(\rho):=\partial_{\rho}\tilde{P}(\rho,\delta)$, we obtain from the definition (1.5):
$\tilde{c}^{2}(\rho)=P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}.$ (2.2)
Lemma 2.1 (Inherited Convexity).
If $P(\rho)$ satisfies conditions (2.1), then for all $\rho>\delta$:
  1. 1.
    $\tilde{c}^{2}(\rho)>0$ (Strict Hyperbolicity),
  2. 2.
    $2\tilde{c}^{2}(\rho)+\rho(\tilde{c}^{2})^{\prime}(\rho)>0$ (Genuine Nonlinearity).
Proof.
1. Let $g(\rho)=\rho^{2}P^{\prime}(\rho)$. Compute $g^{\prime}(\rho)=\rho(2P^{\prime}+\rho P^{\prime\prime})>0$. Since $\rho>\delta$, monotonicity implies $g(\rho)>g(\delta)$, which gives $\rho^{2}\tilde{P}^{\prime}(\rho)=g(\rho)-g(\delta)>0$.
2. We compute the convexity term for the perturbed pressure. Using $\tilde{P}^{\prime}=\tilde{c}^{2}$:
$2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)$ $=2\left(P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}\right)+\rho\frac{d}{d\rho}\left(P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}\right)$
$=2P^{\prime}(\rho)-\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}+\rho P^{\prime\prime}(\rho)+\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}$
$=2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho).$
The singular terms cancel exactly. Since the original pressure satisfies $2P^{\prime}+\rho P^{\prime\prime}>0$, the perturbed pressure satisfies the exact same condition. ∎
2.2 Eigenstructure and Riemann Invariants
Let the vector of conservative variables be $\mathbf{U}=(\rho,m)^{T}$ with momentum $m=\rho u$. The flux vector is given by:
$\mathbf{F}(\mathbf{U})=\begin{pmatrix}m\\ \frac{m^{2}}{\rho}+\tilde{P}(\rho)\end{pmatrix}$. (2.3)
The system in quasi-linear form is $\mathbf{U}_{t}+A(\mathbf{U})\mathbf{U}_{x}=0$, where $A(\mathbf{U})=\nabla_{\mathbf{U}}\mathbf{F}$ is the Jacobian matrix. We compute the partial derivatives of the flux components $F_{1}=m$ and $F_{2}=m^{2}/\rho+\tilde{P}(\rho)$:
$\frac{\partial F_{1}}{\partial\rho}$ $=0,$ $\frac{\partial F_{1}}{\partial m}$ $=1,$
$\frac{\partial F_{2}}{\partial\rho}$ $=-\frac{m^{2}}{\rho^{2}}+\tilde{P}^{\prime}(\rho)=\tilde{c}^{2}-u^{2},$ $\frac{\partial F_{2}}{\partial m}$ $=\frac{2m}{\rho}=2u.$
Thus, the Jacobian matrix is:
$A(\mathbf{U})=\begin{pmatrix}0&1\\ \tilde{c}^{2}-u^{2}&2u\end{pmatrix}$. (2.4)
The eigenvalues $\lambda$ are determined by the characteristic equation $\det(A(\mathbf{U})-\lambda I)=0$:
$\det\begin{pmatrix}-\lambda&1\\ \tilde{c}^{2}-u^{2}&2u-\lambda\end{pmatrix}=-\lambda(2u-\lambda)-(\tilde{c}^{2}-u^{2})=\lambda^{2}-2u\lambda+(u^{2}-\tilde{c}^{2})=0.$ (2.5)
The eigenvalues are:
$\lambda_{1}=u-\tilde{c}(\rho),\quad\lambda_{2}=u+\tilde{c}(\rho).$
The Riemann invariants $z,w$ are quantities constant along integral curves of the eigenvectors. They satisfy $du=\pm\frac{\tilde{c}(\rho)}{\rho}d\rho$. We define:
$z(\rho,u)=u-\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds,\quad w(\rho,u)=u+\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds.$ (2.6)
The integral is well-defined near $\rho=\delta$ because $\tilde{c}(\delta)=0$ and $(\tilde{c}^{2})^{\prime}(\delta)>0$, implying $\tilde{c}(s)\sim\sqrt{s-\delta}$ near the boundary.
Remark 2.2 (Consistency Limit).
It is straightforward to verify that as $\delta\to 0$, the perturbed invariants converge to the classical ones:
$\lim_{\delta\to 0}\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds=\int_{0}^{\rho}\frac{\sqrt{P^{\prime}(s)}}{s}\,ds.$
This ensures that the invariant regions we construct later will converge to bounded regions for the original Euler system.
Proposition 2.3 (Explicit Form for Polytropic Gas).
For $P(\rho)=\kappa\rho^{\gamma}$, the Riemann invariants are:
$z(\rho,u)=u-\sqrt{\kappa\gamma}\int_{\delta}^{\rho}\frac{\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}}{s^{2}}\,ds,\quad w(\rho,u)=u+\sqrt{\kappa\gamma}\int_{\delta}^{\rho}\frac{\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}}{s^{2}}\,ds.$ (2.7)
Proof.
Substituting $P^{\prime}(s)=\kappa\gamma s^{\gamma-1}$ into (2.2) yields $\tilde{c}(s)=\frac{\sqrt{\kappa\gamma}}{s}\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}$. The result follows by integration. ∎
Remark 2.4 (Comparison with Flux Modification).
It is worth noting why the mass flux modification in [4] is unnecessary here. Lu modified the continuity equation to $\rho_{t}+(\rho u-2\delta u)_{x}=0$ to force the eigenvalues to behave as $\lambda\approx u\pm\frac{\rho-2\delta}{\rho}\sqrt{P^{\prime}}$, thereby creating a degenerate boundary at $\rho=2\delta$. In contrast, our approach targets the sound speed directly. By constructing $\tilde{P}$ such that $\tilde{P}^{\prime}(\delta)=0$, we achieve the degeneracy $\lambda\to u$ at $\rho=\delta$ intrinsically. Furthermore, Lemma 2.1 shows that this construction preserves the convexity structure required for entropy analysis, rendering the auxiliary mass perturbation mathematically superfluous.
3 Entropy Analysis
To apply the compensated compactness method, we require a strictly convex entropy pair $(\eta,q)$ that satisfies the entropy identity $\partial_{t}\eta(\mathbf{U})+\partial_{x}q(\mathbf{U})=0$ for smooth solutions.
3.1 Construction of the Mechanical Energy Entropy Pair
The mechanical energy serves as the canonical entropy for the isentropic system. We define the specific internal energy $e(\rho)$ associated with the perturbed pressure via the thermodynamic relation $de=-\tilde{P}d(1/\rho)=\frac{\tilde{P}}{\rho^{2}}d\rho$. Normalizing such that $e(\delta)=0$, we obtain:
$e(\rho)=\int_{\delta}^{\rho}\frac{\tilde{P}(s)}{s^{2}}\,ds.$ (3.1)
The mechanical energy entropy (per unit volume) is defined as $\eta^{*}=\frac{1}{2}\rho u^{2}+\rho e(\rho)$. In terms of the conservative variables $(\rho,m)$ with $m=\rho u$, this reads:
$\eta^{*}(\rho,m)=\frac{m^{2}}{2\rho}+\rho\int_{\delta}^{\rho}\frac{\tilde{P}(s)}{s^{2}}\,ds.$ (3.2)
The corresponding entropy flux $q^{*}$ is given by:
$q^{*}(\rho,m)=\left(\eta^{*}+\tilde{P}(\rho)\right)u=\left(\frac{m^{2}}{2\rho}+\rho e(\rho)+\tilde{P}(\rho)\right)\frac{m}{\rho}.$ (3.3)
Verification of the Entropy Pair.
We verify the compatibility condition by computing the gradients with respect to the conservative variables $\mathbf{U}=(\rho,m)$. First, from the definition of $e(\rho)$, we have $e^{\prime}(\rho)=\frac{\tilde{P}(\rho)}{\rho^{2}}$. It follows that:
$(\rho e(\rho))^{\prime}=e(\rho)+\rho e^{\prime}(\rho)=e(\rho)+\frac{\tilde{P}(\rho)}{\rho}.$
The gradient of the entropy $\eta^{*}$ is computed as:
$\eta^{*}_{\rho}=-\frac{m^{2}}{2\rho^{2}}+(\rho e)^{\prime}=-\frac{1}{2}u^{2}+e(\rho)+\frac{\tilde{P}(\rho)}{\rho},$
$\eta^{*}_{m}=\frac{m}{\rho}=u.$
Substituting these into the entropy equation for smooth solutions:
$\partial_{t}\eta^{*}+\partial_{x}q^{*}$ $=\eta^{*}_{\rho}\rho_{t}+\eta^{*}_{m}m_{t}+\partial_{x}q^{*}$
$=\left(-\frac{1}{2}u^{2}+e+\frac{\tilde{P}}{\rho}\right)(-\partial_{x}m)+u\left(-\partial_{x}(\rho u^{2}+\tilde{P})\right)+\partial_{x}\left((\eta^{*}+\tilde{P})u\right).$
A direct algebraic expansion confirms that the terms sum to zero. Thus, $(\eta^{*},q^{*})$ constitutes a valid entropy-flux pair. ∎
Theorem 3.1 (Strict Convexity).
The entropy $\eta^{*}(\rho,m)$ is strictly convex with respect to the conservative variables for $\rho>\delta$.
Proof.
We compute the Hessian matrix $\nabla^{2}\eta^{*}$:
$\nabla^{2}\eta^{*}=\begin{pmatrix}\eta^{*}_{\rho\rho}&\eta^{*}_{\rho m}\\ \eta^{*}_{m\rho}&\eta^{*}_{mm}\end{pmatrix}$.
The second derivatives are given by:
$\eta^{*}_{mm}=\frac{\partial}{\partial m}\left(\frac{m}{\rho}\right)=\frac{1}{\rho},\quad\eta^{*}_{\rho m}=\frac{\partial}{\partial\rho}\left(\frac{m}{\rho}\right)=-\frac{m}{\rho^{2}}=-\frac{u}{\rho}.$
For $\eta^{*}_{\rho\rho}$, we differentiate $\eta^{*}_{\rho}$ with respect to $\rho$:
$\eta^{*}_{\rho\rho}$ $=\frac{\partial}{\partial\rho}\left(-\frac{u^{2}}{2}+e(\rho)+\frac{\tilde{P}}{\rho}\right)\quad(\text{noting }u=m/\rho)$
$=\frac{\partial}{\partial\rho}\left(-\frac{m^{2}}{2\rho^{2}}\right)+e^{\prime}(\rho)+\frac{\tilde{P}^{\prime}\rho-\tilde{P}}{\rho^{2}}$
$=\frac{m^{2}}{\rho^{3}}+\frac{\tilde{P}}{\rho^{2}}+\frac{\tilde{P}^{\prime}}{\rho}-\frac{\tilde{P}}{\rho^{2}}$
$=\frac{u^{2}}{\rho}+\frac{\tilde{P}^{\prime}(\rho)}{\rho}.$
To check for convexity, we examine the determinant of the Hessian:
$\det(\nabla^{2}\eta^{*})$ $=\eta^{*}_{\rho\rho}\eta^{*}_{mm}-(\eta^{*}_{\rho m})^{2}$
$=\left(\frac{u^{2}}{\rho}+\frac{\tilde{P}^{\prime}}{\rho}\right)\frac{1}{\rho}-\left(-\frac{u}{\rho}\right)^{2}$
$=\frac{u^{2}}{\rho^{2}}+\frac{\tilde{P}^{\prime}(\rho)}{\rho^{2}}-\frac{u^{2}}{\rho^{2}}$
$=\frac{\tilde{P}^{\prime}(\rho)}{\rho^{2}}.$
Since $\rho>\delta$, we have $\rho>0$. By Lemma 2.1, $\tilde{P}^{\prime}(\rho)=\tilde{c}^{2}(\rho)>0$ for $\rho>\delta$. Furthermore, the trace is $\text{tr}(\nabla^{2}\eta^{*})=\frac{u^{2}+\tilde{P}^{\prime}}{\rho}+\frac{1}{\rho}>0$. Positive trace and positive determinant imply that the Hessian is positive definite. Hence, $\eta^{*}$ is strictly convex for $\rho>\delta$. ∎
Remark 3.2 (Degeneracy of Entropy at Cut-off).
While the perturbed pressure $\tilde{P}(\rho,\delta)$ preserves the genuine nonlinearity condition ($2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}>0$) for all $\rho>\delta$, the strict convexity of the mechanical energy entropy $\eta^{*}$ holds only for $\rho>\delta$.
This is a direct consequence of the constructed acoustic degeneracy at the cut-off density: $\tilde{P}^{\prime}(\delta)=0$. Consequently, the determinant of the Hessian vanishes at the boundary:
$\det(\nabla^{2}\eta^{*})\Big|_{\rho=\delta}=0.$
The entropy $\eta^{*}$ therefore exhibits a degenerate convexity (or loss of strict convexity) at the boundary $\rho=\delta$, which geometrically corresponds to the flattening of the entropy surface.
4 Viscous Regularization and A Priori Estimates
We consider the approximate system with artificial viscosity $\varepsilon>0$:
$\left\{\begin{aligned} &\rho_{t}+(\rho u)_{x}=\varepsilon\rho_{xx},\\ &(\rho u)_{t}+(\rho u^{2}+\tilde{P})_{x}=\varepsilon(\rho u)_{xx}.\end{aligned}\right$. (4.1)
4.1 Evolution of Riemann Invariants
Recall $\lambda_{2}=u+\tilde{c}(\rho)$ and $w(\rho,u)=u+\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds$. We compute the transport equation for $w$.
Step 1: Rewrite the momentum equation.
Using
$(\rho u)_{t}+(\rho u^{2})_{x}=\rho(u_{t}+uu_{x})+u(\rho_{t}+(\rho u)_{x}),$
and inserting the mass equation $\rho_{t}+(\rho u)_{x}=\varepsilon\rho_{xx}$ gives
$(\rho u)_{t}+(\rho u^{2})_{x}=\rho(u_{t}+uu_{x})+\varepsilon u\rho_{xx}.$
Expanding the viscous term, $\varepsilon(\rho u)_{xx}=\varepsilon(\rho u_{xx}+2\rho_{x}u_{x}+u\rho_{xx})$. Subtracting the mass equation multiplied by $u$ yields:
$\rho(u_{t}+uu_{x})+\tilde{P}_{x}=\varepsilon\rho u_{xx}+2\varepsilon\rho_{x}u_{x}.$
Dividing by $\rho$ and using $\tilde{P}_{x}=\tilde{c}^{2}\rho_{x}$ gives
$u_{t}+uu_{x}+\frac{\tilde{c}^{2}}{\rho}\rho_{x}=\varepsilon u_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}u_{x}.$ (4.2)
Step 2: Compute $w_{t}+\lambda_{2}w_{x}$.
Using $w_{t}+\lambda_{2}w_{x}=(u_{t}+\lambda_{2}u_{x})+w_{\rho}(\rho_{t}+\lambda_{2}\rho_{x})$ and the mass equation,
$\rho_{t}+\lambda_{2}\rho_{x}=\varepsilon\rho_{xx}-u\rho_{x}-\rho u_{x}+(u+\tilde{c})\rho_{x}.$
Substituting (4.2),
$w_{t}+\lambda_{2}w_{x}$ $=\Big(\varepsilon u_{xx}+\tfrac{2\varepsilon}{\rho}\rho_{x}u_{x}-\tfrac{\tilde{c}^{2}}{\rho}\rho_{x}-uu_{x}+\lambda_{2}u_{x}\Big)$
$\quad+\tfrac{\tilde{c}}{\rho}\big(\varepsilon\rho_{xx}-u\rho_{x}-\rho u_{x}+\lambda_{2}\rho_{x}\big).$
Step 3: Exact cancellation of inviscid terms.
Using $\lambda_{2}=u+\tilde{c}$ and $w_{\rho}=\tilde{c}/\rho$, all non-viscous terms involving $u_{x}$ and $\rho_{x}$ cancel, leaving
$w_{t}+\lambda_{2}w_{x}=\varepsilon u_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}u_{x}+\frac{\varepsilon\tilde{c}}{\rho}\rho_{xx}.$
Step 4: Express $u_{xx}$ via $w_{xx}$.
Since $w_{x}=u_{x}+\frac{\tilde{c}}{\rho}\rho_{x}$,
$w_{xx}=u_{xx}+\frac{\tilde{c}}{\rho}\rho_{xx}+\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}\rho_{x}^{2}.$
Substituting $u_{xx}+\frac{\tilde{c}}{\rho}\rho_{xx}=w_{xx}-(\frac{\tilde{c}}{\rho})^{\prime}\rho_{x}^{2}$ into the evolution equation:
$w_{t}+\lambda_{2}w_{x}=\varepsilon w_{xx}-\varepsilon\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}\rho_{x}^{2}+\frac{2\varepsilon}{\rho}\rho_{x}w_{x}-\frac{2\varepsilon\tilde{c}}{\rho^{2}}\rho_{x}^{2}.$
Step 5: Final evolution equation.
Collecting terms,
$w_{t}+\lambda_{2}w_{x}=\varepsilon w_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}w_{x}-\varepsilon\rho_{x}^{2}\left[\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}+\frac{2\tilde{c}}{\rho^{2}}\right].$ (4.3)
By symmetry, noticing that $z_{xx}=u_{xx}-(\frac{\tilde{c}}{\rho}\rho_{x})_{x}$, the sign of the dissipative term reverses:
$z_{t}+\lambda_{1}z_{x}=\varepsilon z_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}z_{x}+\varepsilon\rho_{x}^{2}\left[\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}+\frac{2\tilde{c}}{\rho^{2}}\right].$ (4.4)
Step 6: Positivity of $K(\rho)$.
Define
$K(\rho)=\Big(\frac{\tilde{c}}{\rho}\Big)^{\prime}+\frac{2\tilde{c}}{\rho^{2}}.$
Since $(\tilde{c}/\rho)^{\prime}=(\rho\tilde{c}^{\prime}-\tilde{c})/\rho^{2}$,
[…]
Using $\tilde{c}^{2}=\tilde{P}^{\prime}$,
$\frac{d}{d\rho}(\rho\tilde{c})=\frac{2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}}{2\tilde{c}}$.
Therefore
$K(\rho)=\frac{2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)}{2\rho^{2}\tilde{c}(\rho)}.$ (4.5)
Since $2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}>0$ for $\rho>\delta$, we obtain $K(\rho)>0$.
4.2 Invariant Regions via the Chueh-Conley-Smoller Theory
Now we establish uniform $L^{\infty}$ bounds by appealing to the general theory of positively invariant regions for systems of nonlinear parabolic equations, as developed by Chueh, Conley, and Smoller [8]. This framework provides a rigorous geometric justification for the bounds in the presence of coupled diffusion and convection.
Consider the viscous system written in terms of the conservative variables $\mathbf{U}=(\rho,m)^{T}$ with $m=\rho u$:
$\mathbf{U}_{t}+\mathbf{F}(\mathbf{U})_{x}=\varepsilon\mathbf{U}_{xx},$ (4.6)
where the flux is $\mathbf{F}(\mathbf{U})=(m,m^{2}/\rho+\tilde{P}(\rho))^{T}$ and the diffusion matrix is the scalar matrix $D=\varepsilon I_{2}$.
We define the phase-space region $\Sigma\subset\mathbb{R}^{2}$ as the intersection of the regions defined by the Riemann invariants and the cut-off density:
$\Sigma=\left\{\mathbf{U}\in\mathbb{R}^{2}:z(\mathbf{U})\geq C_{\inf},\quad w(\mathbf{U})\leq C_{\sup},\quad\rho\geq\delta\right\},$ (4.7)
where $C_{\inf}=\inf_{x}z(\mathbf{U}_{0}(x))$ and $C_{\sup}=\sup_{x}w(\mathbf{U}_{0}(x))$. Note that the third condition $\rho\geq\delta$ is geometrically redundant if the integral $\int_{\delta}^{\rho}\frac{\tilde{c}}{s}ds$ spans the full range, but strictly speaking, it represents the degeneracy boundary $w(\mathbf{U})-z(\mathbf{U})=0$.
Lemma 4.1 (Convexity of Invariant Regions).
Suppose the perturbed pressure satisfies the genuine nonlinearity condition $2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)>0$ for $\rho>\delta$. Then, the phase-space region $\Sigma$ defined by the intersection of the sets:
$\mathcal{R}_{w}=\{(\rho,m):w(\rho,m)\leq C_{\sup}\},\quad\mathcal{R}_{z}=\{(\rho,m):z(\rho,m)\geq C_{\inf}\},\quad\mathcal{R}_{\rho}=\{(\rho,m):\rho\geq\delta\}$
is a convex subset of $\mathbb{R}^{2}$.
Proof.
The region $\mathcal{R}_{\rho}$ is a half-plane $\rho\geq\delta$, which is trivially convex. To prove the convexity of $\mathcal{R}_{w}$ and $\mathcal{R}_{z}$, we examine the curvature of their boundary curves in the $(\rho,m)$-plane.
1. Convexity of $\mathcal{R}_{w}$: The boundary of $\mathcal{R}_{w}$ is the level curve $w(\rho,u)=C_{\sup}$. Expressing $m=\rho u$ as a function of $\rho$ along this curve:
$m_{w}(\rho)=\rho u=\rho\left(C_{\sup}-\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds\right).$
Differentiating with respect to $\rho$, and noting that along this curve $du=-\frac{\tilde{c}}{\rho}d\rho$:
$m_{w}^{\prime}(\rho)=u+\rho u^{\prime}=u+\rho\left(-\frac{\tilde{c}}{\rho}\right)=u-\tilde{c}$.
We compute the second derivative to determine the curvature:
$m_{w}^{\prime\prime}(\rho)=\frac{d}{d\rho}(u-\tilde{c})=u^{\prime}-\tilde{c}^{\prime}(\rho)=-\frac{\tilde{c}}{\rho}-\tilde{c}^{\prime}(\rho)=-\frac{1}{\rho}(\tilde{c}(\rho)+\rho\tilde{c}^{\prime}(\rho)).$
Recall that $\tilde{c}^{2}(\rho)=\tilde{P}^{\prime}(\rho)$. Differentiating gives $2\tilde{c}\tilde{c}^{\prime}=\tilde{P}^{\prime\prime}$. Thus $\tilde{c}^{\prime}=\frac{\tilde{P}^{\prime\prime}}{2\tilde{c}}$. Substituting this back:
$\tilde{c}+\rho\tilde{c}^{\prime}=\tilde{c}+\frac{\rho\tilde{P}^{\prime\prime}}{2\tilde{c}}=\frac{2\tilde{c}^{2}+\rho\tilde{P}^{\prime\prime}}{2\tilde{c}}=\frac{2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}}{2\tilde{c}}$.
Since $\tilde{c}>0$ for $\rho>\delta$ and the genuine nonlinearity condition $2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}>0$ holds (as proved in Lemma 2.1), we have $\tilde{c}+\rho\tilde{c}^{\prime}>0$. Consequently, $m_{w}^{\prime\prime}(\rho)<0$. This implies that the function $m_{w}(\rho)$ is strictly concave. The sublevel set $\mathcal{R}_{w}$ (the region ”below” this concave curve) is therefore a convex set.
2. Convexity of $\mathcal{R}_{z}$: Similarly, the boundary of $\mathcal{R}_{z}$ is $z(\rho,u)=C_{\inf}$, which implies $u=C_{\inf}+\int\frac{\tilde{c}}{s}$. The slope of the boundary curve $m_{z}(\rho)$ is:
$m_{z}^{\prime}(\rho)=u+\rho u^{\prime}=u+\rho\left(\frac{\tilde{c}}{\rho}\right)=u+\tilde{c}$.
The second derivative is:
$m_{z}^{\prime\prime}(\rho)=u^{\prime}+\tilde{c}^{\prime}=\frac{\tilde{c}}{\rho}+\tilde{c}^{\prime}=\frac{1}{\rho}(\tilde{c}+\rho\tilde{c}^{\prime})>0.$
Since $m_{z}^{\prime\prime}(\rho)>0$, the boundary curve is strictly convex. The superlevel set $\mathcal{R}_{z}$ (the region ”above” this convex curve) is therefore a convex set.
Since the intersection of convex sets is convex, $\Sigma=\mathcal{R}_{w}\cap\mathcal{R}_{z}\cap\mathcal{R}_{\rho}$ is a convex region in the conservative phase space. ∎
Theorem 4.2 (Positively Invariant Regions).
Assume the initial data $\mathbf{U}_{0}(x)$ lies in $\Sigma$ for all $x\in\mathbb{R}$. Then, the solution $\mathbf{U}(x,t)$ of the viscous system (4.6) remains in $\Sigma$ for all $t>0$. Consequently, the density satisfies the uniform lower bound $\rho(x,t)\geq\delta$.
Proof.
According to the foundational theory of Chueh, Conley, and Smoller [8] (Theorem 2.1 and Corollary 3.2), a region $\Sigma$ defined by a set of inequalities $\{G_{i}(\mathbf{U})\leq 0\}_{i=1}^{k}$ is positively invariant for the system (4.6) if the following conditions hold on the boundary $\partial\Sigma$:
  1. 1.
    Eigenvector Condition: At any boundary point $\mathbf{U}^{*}$ where $G_{i}(\mathbf{U}^{*})=0$, the gradient $\nabla_{\mathbf{U}}G_{i}(\mathbf{U}^{*})$ is a left eigenvector of the diffusion matrix $D$.
  2. 2.
    Quasi-Convexity: The region $\Sigma$ is convex (or the functions $G_{i}$ are quasi-convex).
  3. 3.
    Flux Condition (Characteristic Boundary): The gradient $\nabla_{\mathbf{U}}G_{i}$ is a left eigenvector of the Flux Jacobian $\mathbf{F}^{\prime}(\mathbf{U})$, or equivalently, the vector field points tangent to the boundary.
We verify these conditions for the defining constraints of $\Sigma$:
1. The Diffusion Condition: Since the artificial viscosity is scalar ($D=\varepsilon I_{2}$), every non-zero vector in $\mathbb{R}^{2}$ is a left eigenvector of $D$. Thus, this condition is trivially satisfied for the gradients of $w$, $z$, and $\rho$ everywhere on the boundary.
2. Convexity of $\Sigma$: The region $\Sigma$ is the intersection of three sets.
  • The constraint $\rho\geq\delta$ defines a convex half-plane in the $(\rho,m)$ coordinates.
[…]
    $m^{\prime\prime}(\rho)=-\frac{2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)}{2\rho\tilde{c}(\rho)}.$
    Since our weighted pressure $\tilde{P}$ inherits the genuine nonlinearity condition $2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}>0$ for all $\rho>\delta$ (Lemma 2.1), we have $m^{\prime\prime}(\rho)<0$. This implies the boundary is concave and the sublevel set $\{w\leq C_{\sup}\}$ is convex. Similarly, for $z=\text{const}$, $m^{\prime\prime}(\rho)>0$, implying convexity of the superlevel set.
  • Thus, $\Sigma$ is the intersection of convex sets and is therefore convex.
3. The Flux Condition on Boundaries: We examine the invariance of the boundaries determined by the hyperbolic character of the system.
  • Upper and Left Boundaries ($w=C_{\sup}$ and $z=C_{\inf}$): By the definition of Riemann invariants, their gradients $\nabla_{\mathbf{U}}w$ and $\nabla_{\mathbf{U}}z$ are the left eigenvectors of the Flux Jacobian $\mathbf{F}^{\prime}(\mathbf{U})$ corresponding to eigenvalues $\lambda_{2}=u+\tilde{c}$ and $\lambda_{1}=u-\tilde{c}$, respectively. Thus, these boundaries are characteristic surfaces. The convective flux is tangent to the boundary $\partial\Sigma$, preventing the solution from crossing it.
  • The Lower Boundary ($\rho=\delta$): This boundary corresponds to the degeneracy locus $w(\mathbf{U})=z(\mathbf{U})$. We observe that as $\rho\to\delta$, the sound speed vanishes ($\tilde{c}\to 0$). Consequently, the two eigenvalues coalesce:
    $\lim_{\rho\to\delta}\lambda_{1}=\lim_{\rho\to\delta}\lambda_{2}=u.$
    Correspondingly, the gradients of the Riemann invariants become collinear. Since both $\nabla w$ and $\nabla z$ are left eigenvectors corresponding to the limit eigenvalue $u$, any linear combination of them is also a left eigenvector for $u$. Specifically, the gradient of the boundary defining function G=zwG=z-w satisfies:
    $\nabla_{\mathbf{U}}(z-w)\cdot\mathbf{F}^{\prime}(\mathbf{U})\Big|_{\rho=\delta}=(\nabla z\cdot\mathbf{F}^{\prime}-\nabla w\cdot\mathbf{F}^{\prime})=(u\nabla z-u\nabla w)=u\nabla_{\mathbf{U}}(z-w).$
    This proves that the boundary $\rho=\delta$ is a characteristic boundary with characteristic speed $u$. The flow is strictly tangential to this boundary in the phase space. Physically, this means the fluid velocity matches the boundary propagation speed, making it impossible for the particle trajectories to cross $\rho=\delta$.
Since all geometric conditions of the CCS theory are satisfied, $\Sigma$ is a positively invariant region. This establishes the uniform bounds independent of $\varepsilon$:
$\delta\leq\rho^{\varepsilon}(x,t)\leq M,\quad|u^{\varepsilon}(x,t)|\leq M.$
4.3 A Geometric Re-interpretation of the work of Lu
In this section, we provide a geometric verification of the invariant regions constructed in the seminal work of Lu [4]. While Lu originally established the existence of these regions via coordinate transformations and energy estimates, we demonstrate here that his result can be derived directly from the Chueh-Conley-Smoller (CCS) theory [8]. This analysis reveals that the modification of the mass flux in [4] serves precisely to satisfy the tangential flow condition required by the CCS framework.
4.3.1 Lu’s Modified System
Recall the regularization system proposed in [4]:
$\left\{\begin{aligned} &\rho_{t}+(\rho u-2\delta u)_{x}=\varepsilon\rho_{xx},\\ &(\rho u)_{t}+(\rho u^{2}-\delta u^{2}+P_{1}(\rho,\delta))_{x}=\varepsilon(\rho u)_{xx}.\end{aligned}\right$. (4.8)
The perturbed pressure is defined as $P_{1}(\rho,\delta)=\int_{2\delta}^{\rho}\frac{s-2\delta}{s}P^{\prime}(s)\,ds$. The invariant region proposed by Lu is:
$\Sigma^{Lu}=\left\{(\rho,u):\rho\geq 2\delta,\quad z^{Lu}\geq C_{1},\quad w^{Lu}\leq C_{2}\right\},$ (4.9)
where $z^{Lu},w^{Lu}$ are the Riemann invariants associated with the system (4.8).
4.3.2 Proof via CCS Theory
We now prove the invariance of $\Sigma^{Lu}$ using the geometric criteria of [8].
Theorem 4.3.
The region $\Sigma^{Lu}$ is positively invariant for the system (4.8) with respect to the CCS theory.
Proof.
Let $\mathbf{U}=(\rho,m)^{T}$. The system (4.8) takes the form $\mathbf{U}_{t}+\mathbf{F}_{Lu}(\mathbf{U})_{x}=\varepsilon\mathbf{U}_{xx}$. The diffusion matrix is scalar ($D=\varepsilon I$), satisfying the diffusion condition trivially. We focus on the Flux Condition at the boundaries of $\Sigma^{Lu}$.
1. The Hyperbolic Boundaries ($z^{Lu}=C_{1},w^{Lu}=C_{2}$): The system (4.8) is strictly hyperbolic for $\rho>2\delta$. By definition, the gradients of the Riemann invariants are the left eigenvectors of the Jacobian $\mathbf{F}_{Lu}^{\prime}(\mathbf{U})$. Thus, the convective vector field is tangential to the level sets of $z^{Lu}$ and $w^{Lu}$. The standard CCS condition for hyperbolic boundaries is satisfied.
2. The Artificial Boundary ($\rho=2\delta$): The core difficulty lies in the lower boundary $\rho=2\delta$. We must verify that the vector field $\mathbf{F}_{Lu}$ does not point outward across the surface $\rho=2\delta$. Let us examine the eigenvalues of Lu’s system. The modified flux Jacobian has eigenvalues (see Eq. (2.1) in [4]):
$\lambda_{1,2}^{Lu}=u\mp\frac{\rho-2\delta}{\rho}\sqrt{P^{\prime}(\rho)}.$ (4.10)
The crucial observation is the behavior at the boundary $\rho=2\delta$. Substituting $\rho=2\delta$ into (4.10):
$\lambda_{1}^{Lu}=\lambda_{2}^{Lu}=u.$ (4.11)
Geometrically, this implies that the characteristic curves become exactly tangent to the boundary $\rho=2\delta$ in the $(x,t)$ plane. In the phase space $(\rho,m)$, the boundary $\rho=2\delta$ corresponds to the degeneracy locus where the two families of characteristics coincide. Specifically, let $G(\mathbf{U})=2\delta-\rho\leq 0$. The normal vector is $\mathbf{n}=(-1,0)^{T}$. The condition that the flow is tangential means that the boundary propagates at the characteristic speed. Since $\lambda=u$, the fluid particles at the boundary move exactly with the boundary. To see this rigorously using the CCS condition, we verify that $\nabla\rho=(1,0)$ is a left eigenvector of $\mathbf{F}_{Lu}^{\prime}$ at $\rho=2\delta$ corresponding to the eigenvalue $u$. The mass conservation equation in (4.8) can be rewritten as:
$\rho_{t}+u\rho_{x}+(\rho-2\delta)u_{x}=\varepsilon\rho_{xx}.$
At the boundary where $\rho=2\delta$, the term $(\rho-2\delta)u_{x}$ vanishes. The equation degenerates to a pure convection-diffusion equation:
$\rho_{t}+u\rho_{x}=\varepsilon\rho_{xx}.$
This confirms that the convective transport velocity normal to the boundary $\rho=2\delta$ is exactly $u$. Since the boundary moves with the fluid, there is no cross-boundary flux. Thus, the condition $\nabla G\cdot\mathbf{F}_{Lu}^{\prime}\propto\nabla G$ holds, satisfying the CCS tangency requirement.
Therefore, the boundaries are either characteristic surfaces or degenerate characteristic surfaces where the flow is tangential, and the convexity of the region is guaranteed by the structure of gas dynamics, the region $\Sigma^{Lu}$ is positively invariant. ∎
Remark 4.4 (Geometric Unity with Lu’s Regularization).
It is illuminating to interpret the regularization strategy of Lu [4] and the current weighted pressure approach through the lens of the Chueh-Conley-Smoller theory. Although the two methods employ distinct mechanisms—Lu modifies the mass flux to $\rho u-2\delta u$, whereas we perturb the pressure to $\tilde{P}$—they share the identical geometric essence.
From the perspective of invariant regions, the core requirement is that the boundary of the domain must be a characteristic surface where the vector field is tangential (the Flux Condition).
  • Lu’s Approach: The flux modification in [4] is algebraically designed to force the eigenvalues to degenerate to $\lambda=u$ at the boundary $\rho=2\delta$, thereby satisfying the tangential condition via transport retardation.
  • Our Approach: We achieve the same characteristic degeneracy $\lambda=u\pm\tilde{c}\to u$ at $\rho=\delta$ by geometrically flattening the pressure such that the sound speed vanishes ($\tilde{c}\to 0$).
Thus, both methods represent different pathways to satisfy the same geometric constraint required for the existence of invariant regions. However, our approach distinguishes itself by satisfying this geometric condition while strictly preserving the physical law of mass conservation.
4.4 Energy Dissipation and $H^{-1}$ Compactness
We now establish compactness of entropy dissipation measures. Let $(\eta^{*},q^{*})$ be the strictly convex mechanical entropy–flux pair. For viscous solutions of (4.1), direct computation gives the entropy identity:
$\partial_{t}\eta^{*}(\mathbf{U}^{\varepsilon})+\partial_{x}q^{*}(\mathbf{U}^{\varepsilon})=\varepsilon\eta^{*}_{xx}-\varepsilon(\mathbf{U}^{\varepsilon}_{x})^{T}\nabla^{2}\eta^{*}(\mathbf{U}^{\varepsilon})\mathbf{U}^{\varepsilon}_{x},$ (4.12)
where $\mathbf{U}^{\varepsilon}=(\rho^{\varepsilon},\rho^{\varepsilon}u^{\varepsilon})^{T}$.
Dissipation structure. Using $m_{x}=\rho_{x}u+\rho u_{x}$ and the Hessian derived in Section 3, one obtains the dissipation term:
$D^{\varepsilon}:=\varepsilon(\mathbf{U}^{\varepsilon}_{x})^{T}\nabla^{2}\eta^{*}\mathbf{U}^{\varepsilon}_{x}=\varepsilon\Big(\frac{\tilde{P}^{\prime}(\rho^{\varepsilon})}{\rho^{\varepsilon}}(\rho^{\varepsilon}_{x})^{2}+\rho^{\varepsilon}(u^{\varepsilon}_{x})^{2}\Big)\geq 0.$
Integrating (4.12) gives uniform bounds on the dissipation, implying $\sqrt{\varepsilon}\rho_{x}^{\varepsilon}$ and $\sqrt{\varepsilon}u_{x}^{\varepsilon}$ are uniformly bounded in $L^{2}_{\text{loc}}$.
Compactness of entropy production. Decompose the entropy production as:
$\partial_{t}\eta(\mathbf{U}^{\varepsilon})+\partial_{x}q(\mathbf{U}^{\varepsilon})=I_{1}^{\varepsilon}+I_{2}^{\varepsilon},$
where $I_{1}^{\varepsilon}=\varepsilon\eta_{xx}$ and $I_{2}^{\varepsilon}=-D^{\varepsilon}$.
  1. 1.
    Since $\sqrt{\varepsilon}\mathbf{U}_{x}^{\varepsilon}$ is bounded in $L^{2}$, $\varepsilon\eta_{x}\to 0$ in $L^{2}$. Thus $I_{1}^{\varepsilon}\to 0$ in $H^{-1}_{\mathrm{loc}}$.
  2. 2.
    The term $I_{2}^{\varepsilon}$ is bounded in $L^{1}$, hence bounded in the space of measures $\mathcal{M}$.
Applying Murat’s Lemma [7], the divergence of the entropy pair lies in a compact subset of $H^{-1}_{\mathrm{loc}}$. This allows the application of the compensated compactness theory to extract a convergent subsequence.
5 Global Existence of Weak Solutions for Fixed $\delta>0$
In this section, we pass to the limit as $\varepsilon\to 0$ to establish the global existence of weak entropy solutions for the system with weighted pressure $\tilde{P}$.
Definition 5.1 (Weak Entropy Solution).
A pair of bounded measurable functions $(\rho,u)$ with $\rho\geq\delta$ is called a weak entropy solution to the Cauchy problem for the perturbed system if:
  1. 1.
    It satisfies the equations in the sense of distributions:
    $\iint_{\mathbb{R}\times\mathbb{R}^{+}}\left(\rho\phi_{t}+\rho u\phi_{x}\right)dxdt+\int_{\mathbb{R}}\rho_{0}(x)\phi(x,0)dx=0,$ (5.1)
    $\iint_{\mathbb{R}\times\mathbb{R}^{+}}\left(\rho u\psi_{t}+(\rho u^{2}+\tilde{P}(\rho))\psi_{x}\right)dxdt+\int_{\mathbb{R}}\rho_{0}(x)u_{0}(x)\psi(x,0)dx=0,$ (5.2)
    for all test functions $\phi,\psi\in C_{0}^{\infty}(\mathbb{R}\times[0,\infty))$.
  2. 2.
    It satisfies the entropy inequality:
    $\partial_{t}\eta(\rho,u)+\partial_{x}q(\rho,u)\leq 0$ (5.3)
    in the sense of distributions for any convex entropy pair $(\eta,q)$ where $\eta$ is convex with respect to conservative variables $(\rho,m)$.
Theorem 5.2 (Global Existence).
Let the initial data $(\rho_{0},u_{0})$ be bounded and satisfy $(\rho_{0}(x),u_{0}(x))\in\Sigma$. Then, for any fixed $\delta>0$, there exists a subsequence of approximate solutions $(\rho^{\varepsilon},u^{\varepsilon})$ constructed in Section 4 that converges pointwise almost everywhere to a global weak entropy solution $(\rho,u)$ of the perturbed Euler system as $\varepsilon\to 0$.
Proof.
The proof relies on the compensated compactness method of Tartar [2] and Murat [7].
Step 1: Young Measures. From the a priori estimates in Theorem 4.2, the sequence of viscous solutions $\mathbf{U}^{\varepsilon}=(\rho^{\varepsilon},u^{\varepsilon})$ is uniformly bounded in $L^{\infty}(\mathbb{R}\times\mathbb{R}^{+})$. According to the Tartar-DiPerna theory, there exists a subsequence (still denoted by $\mathbf{U}^{\varepsilon}$) and a family of probability measures $\nu_{x,t}$ (Young measures) supported on the invariant region $\Sigma$, such that for any continuous function $f(\mathbf{U})$,
[…]
Step 2: Div-Curl Lemma Application. In Section 4.4, we established that for any smooth entropy pair $(\eta,q)$, the dissipation measure $\partial_{t}\eta(\mathbf{U}^{\varepsilon})+\partial_{x}q(\mathbf{U}^{\varepsilon})$ is compact in $H^{-1}_{\mathrm{loc}}$. Let $(\eta_{1},q_{1})$ and $(\eta_{2},q_{2})$ be two arbitrary weak entropy pairs. By the Div-Curl Lemma, we have the commutation identity:
$\overline{\eta_{1}q_{2}-\eta_{2}q_{1}}=\bar{\eta}_{1}\bar{q}_{2}-\bar{\eta}_{2}\bar{q}_{1}.$ (5.4)
In terms of the Young measure, this reads:
$\langle\nu,\eta_{1}q_{2}-\eta_{2}q_{1}\rangle=\langle\nu,\eta_{1}\rangle\langle\nu,q_{2}\rangle-\langle\nu,\eta_{2}\rangle\langle\nu,q_{1}\rangle.$ (5.5)
Step 3: Reduction of the Young Measure. To prove strong convergence, we must show that the Young measure $\nu_{x,t}$ reduces to a Dirac mass $\delta_{\mathbf{U}(x,t)}$ at almost every point. We utilize the specific structure of the isentropic Euler equations. As shown in Lemma 2.1, the perturbed pressure $\tilde{P}$ satisfies the genuine nonlinearity condition $2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)>0$ for all $\rho\geq\delta$.
According to the theory developed by DiPerna [3] and Chen [5], for systems satisfying strictly hyperbolic and genuinely nonlinear conditions, the commutation identity (5.5) satisfied by a sufficiently large class of weak entropies implies that the support of $\nu$ is a single point. Specifically, since our perturbed system retains the isentropic structure with a convex pressure law $\tilde{P}$, the Lax entropy pairs generated by the fundamental solution of the Euler-Poisson-Darboux equation can be applied. The non-degeneracy of the mapping from the state space to the entropy space ensures that $\nu_{x,t}$ is a Dirac mass.
Thus, we conclude that $\mathbf{U}^{\varepsilon}(x,t)\to\mathbf{U}(x,t)$ strongly in $L^{1}_{\mathrm{loc}}$.
Step 4: Satisfaction of Equations. Since the convergence is strong, we can pass to the limit in the weak formulation of the viscous system (4.1). The viscous terms vanish as $\varepsilon\to 0$ because $\sqrt{\varepsilon}\mathbf{U}_{x}^{\varepsilon}$ is bounded in $L^{2}$. Therefore, the limit function $\mathbf{U}$ satisfies the Euler equations with pressure $\tilde{P}$ in the sense of distributions. ∎
Acknowledgements
The author is grateful to Professor Yunguang Lu for valuable discussions and insightful guidance related to the global existence theory of isentropic gas dynamics.
Appendix A Derivation and Motivation of the Weighted Pressure Perturbation
In this appendix, we provide the heuristic motivation behind the construction of the weighted pressure perturbation $\tilde{P}(\rho,\delta)$. The specific form (1.5) is not arbitrary; rather, it is derived via a reverse-engineering process designed to enforce boundary degeneracy while strictly preserving the convexity structure of the original equation of state.
A.1 The Structural Invariant
Our primary goal is to construct a perturbed pressure $\tilde{P}$ that satisfies the genuine nonlinearity condition:
$2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)>0,\quad\text{for }\rho>\delta.$ (A.1)
Let us analyze the differential structure of this condition. Observe the identity:
$\frac{d}{d\rho}\left(\rho^{2}P^{\prime}(\rho)\right)=2\rho P^{\prime}(\rho)+\rho^{2}P^{\prime\prime}(\rho)=\rho\left(2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho)\right).$ (A.2)
Define the auxiliary structural function $g(\rho):=\rho^{2}P^{\prime}(\rho)$. The above identity implies that the genuine nonlinearity of the system is equivalent to the strict monotonicity of $g(\rho)$:
$2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho)>0\iff g^{\prime}(\rho)>0.$ (A.3)
A.2 The Shift Strategy
To regularize the system, we seek a new pressure $\tilde{P}$ such that the acoustic speed vanishes at the cut-off density $\delta$, i.e., $\tilde{P}^{\prime}(\delta)=0$, while maintaining the same convexity class as $P$. In terms of the structural function $\tilde{g}(\rho):=\rho^{2}\tilde{P}^{\prime}(\rho)$, the boundary condition becomes:
$\tilde{g}(\delta)=\delta^{2}\tilde{P}^{\prime}(\delta)=0.$ (A.4)
To preserve the convexity property, we impose the strongest possible structural similarity: we require the derivative of the structural function to remain invariant:
$\tilde{g}^{\prime}(\rho)=g^{\prime}(\rho).$ (A.5)
Integrating this condition from $\delta$ to $\rho$ with the initial condition $\tilde{g}(\delta)=0$ yields a simple shift of the original function:
$\tilde{g}(\rho)=g(\rho)-g(\delta).$ (A.6)
Substituting back the definitions of $g$ and $\tilde{g}$, we obtain:
$\rho^{2}\tilde{P}^{\prime}(\rho)=\rho^{2}P^{\prime}(\rho)-\delta^{2}P^{\prime}(\delta).$ (A.7)
A.3 Reconstruction and Exact Cancellation
Solving for $\tilde{P}^{\prime}(\rho)$, we get the derivative of the perturbed pressure:
$\tilde{P}^{\prime}(\rho)=P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}.$ (A.8)
Integrating (A.8) with respect to $\rho$ gives the explicit formula used in this paper:
$\tilde{P}(\rho)=\int_{\delta}^{\rho}\left(P^{\prime}(s)-\frac{\delta^{2}P^{\prime}(\delta)}{s^{2}}\right)\,ds=\int_{\delta}^{\rho}\frac{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}{s^{2}}\,ds.$ (A.9)
Finally, we demonstrate why this specific construction leads to the exact preservation of genuine nonlinearity. Calculating the convexity term for $\tilde{P}$ using (A.8):
$2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)$ $=2\left(P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}\right)+\rho\frac{d}{d\rho}\left(P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}\right)$
$=2P^{\prime}(\rho)-\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}+\rho P^{\prime\prime}(\rho)+\rho\left(\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{3}}\right)$
$=(2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho))\underbrace{-\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}+\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}}_{\text{Exact Cancellation}}.$ (A.10)
Thus, the singular terms introduced by the regularization cancel exactly, leaving the convexity condition determined solely by the original pressure law. This algebraic property is the cornerstone of our global existence proof.